### v1

# Large Language Models Are Cross-Lingual Knowledge-Free Reasoners

Abstract
Large Language Models have demonstrated impressive reasoning capabilities across multiple languages. However, the relationship between capabilities in different languages is less explored. In this work, we decompose the process of reasoning tasks into two separated parts: knowledge retrieval and knowledge-free reasoning, and analyze the cross-lingual transferability of them. With adapted and constructed knowledge-free reasoning datasets, we show that the knowledge-free reasoning capability can be nearly perfectly transferred across various source-target language directions despite the secondary impact of resource in some specific target languages, while cross-lingual knowledge retrieval significantly hinders the transfer. Moreover, by analyzing the hidden states and feed-forward network neuron activation during the reasoning tasks, we show that higher similarity of hidden representations and larger overlap of activated neurons could explain the better cross-lingual transferability of knowledge-free reasoning than knowledge retrieval. Thus, we hypothesize that knowledge-free reasoning embeds in some language-shared mechanism, while knowledge is stored separately in different languages.  [^1] Our code and data is available at: https://github.com/NJUNLP/Knowledge-Free-Reasoning.
Large Language Models Are Cross-Lingual Knowledge-Free Reasoners
[…]
1 Introduction
Large language models (LLMs) today have shown strong multitask and multilingual performance in various domains Huang and Chang (2022), including robust reasoning capabilities across multiple languages Ranaldi and Zanzotto (2023), even for low-resource languages in the training corpus Shi et al. (2022).
Previous study reveals that these multilingual LLMs possess certain ability of multilingual transfer Ye et al. (2023), which means the skills or knowledge learned with one language can be automatically transferred to another language without extra training. However, the effect of such cross-lingual transfer varies across tasks. In knowledge-related tasks, current LLMs show unsatisfactory cross-lingual transfer Gao et al. (2024), while in certain reasoning tasks, more effective transfer is observed Qi et al. (2023). Previous study lacks the analysis on the difference between these tasks, and does not dig further into the specific factors affecting the transfer effectiveness.
In this study, we divide a reasoning task into two separated components: knowledge retrieval and knowledge-free reasoning, and investigate the transfer effectiveness of them in current LLMs. The former means recalling information encountered during training, while the latter refers to organizing the given knowledge in the context into a final answer.
Based on such definition, we perform evaluation and interpretability analysis on LLaMA-2 models Touvron et al. (2023) to compare the cross-lingual transferability of knowledge retrieval and knowledge-free reasoning. In the evaluation part, by controlling the requirement of knowledge retrieval in the StrategyQA (Geva et al., 2021) dataset and constructing a knowledge-free reasoning dataset, we compare the cross-lingual transfer effectiveness of knowledge-related and knowledge-free reasoning. In the interpretability analysis part, we assess the cross-lingual computational similarity of hidden states and Feed-Forward Network (FFN) neural activation to trace the computational process of cross-lingual transfer of knowledge-related and knowledge-free reasoning in LLMs. Our main findings are:
  • Knowledge retrieval significantly hinders cross-lingual transfer. The more knowledge retrieval is required in the task, the lower effectiveness of cross-lingual transfer is observed.
  • The knowledge-free reasoning ability can be effectively transferred to other languages after finetuning in one, while the model’s ability in these languages are less important.
  • The overall cross-lingual computational similarity of knowledge-free reasoning is significantly higher than knowledge-related reasoning. Also, the difference in such similarity level is the most significant in the middle-high layers of the model. This suggests a language-shared mechanism of knowledge-free reasoning in multilingual LLMs.
2 Methods
The work in this paper consists of two parts: evaluation and interpretability analysis. For each part, we employ specific datasets and design certain metrics or measurements.
2.1 Datasets
This part introduces the datasets used in this paper for the analysis of the cross-lingual transfer of reasoning tasks. For datasets available only in English or part of the tested languages, we translate them into other languages with Google Translate.
2.1.1 Adapted StrategyQA dataset
The StrategyQA dataset is a multi-domain reasoning dataset including common sense, arithmetic, symbolic, and logical reasoning, where model is given at least two pieces of evidence from Wikipedia, and required to give "yes" or "no" answers to the questions. An example can be found in Table 3 in Appendix B.
Since Wikipedia is a commonly used source in LLM pretraining, the model can either collect the required information from evidence in the context, or retrieve it from its internal knowledge strorage. Thus, by masking out some of the evidence, we can control the need of knowledge retrieval in the reasoning process, which allows us to analyze the impact of knowledge retrieval on the cross-lingual transfer of reasoning skills. Namely, we design two kinds of scenarios in the experiments:
  • No Fact (NF): The model is given only the questions and must retrieve the necessary knowledge (evidence).
  • With Fact (WF): The model is provided with the questions and some of the evidence. To control the degree of knowledge retrieval needed, we further devide the WF-1, WF-2 and WF-all settings, where one piece, two pieces, and all pieces of evidence is provided for each question.
2.1.2 Knowledge-free reasoning dataset
Although the need for knowledge retrieval can be controlled with the adapted StrategyQA dataset, the questions themselves may still contain implicit requirement for knowledge retrieval. To rule out such effect and focus on the knowledge-free reasoning process, we construct a data that minimizes the need for knowledge retrieval. Inspired by Wei et al. (2022)’s taxonomy of reasoning tasks, we developed the Knowledge-Free Reasoning Dataset (KFRD), which consists of three fundamental reasoning tasks: arithmetic reasoning, symbolic reasoning, and logic reasoning in the 4-choice format. The dataset is constructed by filling GPT-4 Achiam et al. (2023) generated entities or simple English words into a unified template with one input part, one transformation rule, and one options part. Then, using the translated templates, entities and words, the dataset is expanded to all the tested languages. The template and examples can be found in Figure 11 and Table 1 in Appendix A.
Arithmetic Reasoning:
transforming one or two input numbers through mathematical operations into one or two output numbers. The mathematical operations include: basic arithmetic operations (addition, subtraction, multiplication, division), equality, geometric progression, arithmetic progression, and sorting.
Symbolic Reasoning:
transforming 3-5 words from the respective language input through symbolic operations to obtain the output. The symbolic operations include: repetition, addition, deletion, reordering, and their combinations, with up to three operations applied in combination.
Logical Reasoning:
selecting propositions derived from eight given premises using provided logical rules. The logical rules include: Implication Elimination, Conjunction Introduction, Conjunction Elimination, Disjunction Introduction, Disjunction Elimination, and Proof by Contradiction.
2.1.3 Knowledge-related reasoning dataset
To further compare the cross-lingual transfer of knowledge retrieval and knowledge-free reasoning, three knowledge-related reasoning datasets are used in the interpretability analysis. The three datasets are:
MKQA:
The MKQA dataset Longpre et al. (2021) is an open-domain question-answering evaluation set designed to provide a challenging benchmark for multilingual question-answering. The dataset consists of 10,000 question-answer pairs and covers 26 different languages.
BoolQ:
BoolQ (Clark et al., 2019) is a question answering dataset for yes/no questions, comprising 15,942 examples. The questions originate from queries to the Google search engine with Wikipedia pages as answers.
AmbigQA:
AmbigQA (Min et al., 2020) is an open-domain dataset for question answering that includes 14,042 questions. It addresses the issue of ambiguous answers by rephrasing the related questions.
Some examples of these datasets are shown in Table 3 in Appendix B.
2.2 Evaluation metric
In order to measure the model’s cross-lingual transferability while mitigating the difficulty difference in of the datasets, we calculate the Cross-lingual Transfer Ratio (XLTR) following the prior work Gao et al. (2024). The formula is as follows:
$\text{XLTR}(s,t)=(\frac{|C_{s}\cap C_{t}|}{|C_{s}|}-A_{r})/(1-A_{r})$
where $s$ and $t$ denote the source and target languages in the transfer. $C_{x}$ represents the set of correct answers in language $x$, and $A_{r}$ is the accuracy of random choices for the given task. If after training on a task in one specific language (namely, English), the model shows an XLTR score close to 100% in other untrained languages, we say it achieves Fully Cross-Lingual Transfer (FCLT) in this task.
2.3 Interpretability measurements
We employ two metrics to evaluate the cross-lingual computational consistency of the LLM during knowledge-related and knowledge-free reasoning processes.
2.3.1 Cosine similarity of hidden states (CS)
We measure the cosine similarity of the hidden representations across multiple languages during the reasoning process of a same question, in order to observe how the semantic space in the tested languages approximate each other. The similarity is calculated by:
$\text{CS}(x)=\frac{\sum_{n=1}^{N}\sum_{\begin{subarray}{c}a,b\in\mathcal{L},a\neq b\end{subarray}}\frac{\mathbf{h}_{n}^{a}(x)\cdot\mathbf{h}_{n}^{b}(x)}{\left|\mathbf{h}_{n}^{a}(x)\right|\cdot\left|\mathbf{h}_{n}^{b}(x)\right|}}{|\mathcal{L}|(|\mathcal{L}|-1)N}$
where $x$ is a certain question sample, $N$ is the total number of model layers, $\mathcal{L}$ denotes the set of all tested languages, and $\mathbf{h}_{n}^{a}(x)$ is the output hidden states of the $n$-th layer for sample $x$ in language $a$. After that, the cosine similarity of all tested samples are averaged to report the final score.
2.3.2 Neuron Activation Overlap (NAO)
The Neuron Activation Overlap refers to the extent of overlap between activated neurons across multiple languages for the same corpus. The calculation steps are as follow:
  1. 1.
    Input a question in multiple languages into the model and extract the activation values of the last token in the question in each language.
  2. 2.
    Identify activated neurons. A neuron is considered activated if the absolute value of its activation exceeds a certain threshold.
  3. 3.
    Calculate the overlap of activated neurons in all the tested languages.
For a question sample $x$, the formula is as follows:
$\text{NAO}(x)=\frac{\left|\mathcal{L}\right|\cdot\left|\bigcap_{l\in\mathcal{L}}S^{l}(x)\right|}{\sum_{l\in\mathcal{L}}\left|S^{l}(x)\right|}$
where $\mathcal{L}$ is set of languages, and $S^{l}(x)$ is the set of activated neurons on sample $x$ in language $l$. After that, the NAO of all tested samples are averaged to report the final score.
3 Experiment Settings
3.1 Language and model choice
Language choice
To capture a broad spectrum of linguistic diversity, we selected ten languages based on their extensive use and representation of diverse linguistic families, following Gao et al. (2024). The languages selected are English (en), German (de), French (fr), Italian (it), Russian (ru), Polish (pl), Arabic (ar), Hebrew (he), Chinese (zh), and Japanese (ja). Further details are provided in Appendix C.
Model choice
We selected several LLMs including LLaMA-2-7B-Chat (Touvron et al., 2023), BLOOMZ-MT-7B (Muennighoff et al., 2023), Mistral-7B-Instruct-v0.1 (Jiang et al., 2023), and Qwen-1.5-7B-Chat (Bai et al., 2023) to conduct transferability experiments on existing and synthetic datasets.
For experiments focusing on the impact of the training language (Section 4.3.1) and interpretability (Section 4.4), we used LLaMA-2-7B-Chat as a representative model for analysis.
For the experiments evaluating impact of the target language proficiency in specific directions (Section 4.3.2), we selected derived models of LLaMA-2-7B and Mistral-7B to conduct experiments on Arabic and Hebrew respectively, both of which are low-resource languages.
3.2 Fine-tuning settings
We perform LoRA fine-tuning (Hu et al., 2021) on all model blocks in all experiments due to the limited computational resources. More details about fine-tuning can be found in Appendix E.
3.3 Decoding settings
In all experiments, we perform constrained decoding to prevent the model from generating tokens other than the desired choices (Yes/No for StrategyQA, A/B/C/D for KFRD).
In the interpretability analysis, we use the last token of the question to collect the hidden states and neural activation values, because the last input token is used to predict the next token, it gradually incorporates the primary information of the entire sentence, reflecting the overall thought process for the entire problem (Meng et al., 2022; Stolfo et al., 2023; Wu et al., 2024). By focusing on the model’s computational pathway for reasoning rather than calculating the similarity between multilingual sentences, we can better understand how the model processes complex queries. Calculating with an output token, on the other hand, would make it difficult to interpret the reasoning process. Additionally, token counts differ across languages, complicating direct comparisons. Therefore, using the last input token helps in standardizing the analysis across different languages.
To ensure consistency in the final token across different datasets, we added a language-specific ’?’ to certain datasets. For details on these modifications and metrics, please refer to Appendix D.
4 Results
Figure 1: XLTR of different models on StrategyQA. Solid lines: WF results; Dashed lines: NF results. The label of training language (en) is capitalized.
Figure 2: XLTR of LLaMA-2-7B-Chat on StrategyQA under different settings. NF stands for the No Facts setting, while WF-1 and WF-2 indicate that one or two facts are given for each question, respectively. WF-all signifies that all facts are provided.
Figure 3: XLTR on the different parts of KFRD
Figure 4: Accuracy of LLaMA-2-7B-Chat from different languages on three parts of KFRD
Figure 5: XLTR of LLaMA-2-7B-Chat from different languages on three parts of KFRD
4.1 Impact of knowledge retrieval on cross-lingual transfer
We analyze the impact of the amount of retrieved knowledge on cross-lingual transfer in different settings of the StrategyQA dataset. The results for cross-lingual transfer ratio are presented in Figure 1, while the accuracy results are detailed in Figure 12 in Appendix G.
Knowledge retrieval requirement harms cross-lingual transfer
The experimental results indicate that, for all languages, the cross-lingual transfer ratio of all models are significantly higher when the necessary knowledge for reasoning is provided compared to when it is not. This suggests that the requirement for knowledge retrieval significantly hinders the model’s cross-lingual transferability when solving reasoning problems.
Less knowledge retrieval makes better cross-lingual transfer
Additionally, we conduct a more detailed evaluation using the LLaMA-2-7B-Chat model to observe the changes in cross-lingual transfer ratio with varying amounts of knowledge entries. As shown in Figure 2, the experimental results demonstrate that the transfer rate increases as the number of provided knowledge entries increases from none to all required knowledge. This further validates the conclusion that the retrieval of more knowledge significantly impacts cross-lingual transferability.
4.2 The cross-lingual transfer of knowledge-free reasoning
We assess the cross-lingual transferability of the model’s knowledge-free reasoning capabilities. We evaluate the performance of the KFRD by training in English and transferring to the other 9 languages. The resulting cross-lingual transfer ratio are shown in Figure 3, and accuracy results are shown in Figure 13 in Appendix G.
The results demonstrate that the KFRD exhibits extremely high cross-lingual transfer performance for most language pairs. For 7 out of the 9 languages, it can be observed that the cross-lingual transfer performance in knowledge-free reasoning tasks often exceeds 90%, with some instances approaching 100%, thereby achieving near FCLT.
However, for the two lowest-resource languages Arabic and Hebrew, the cross-lingual transferability is significantly poorer. We hypothesize that this may be due to the model’s weaker language proficiency in these two languages, which adversely affects its transferability. Further analysis on this issue is provided in the next section.
4.3 Impact of language proficiency on cross-lingual transfer
In this section, we explore the impact of language proficiency of the training and target language with additional experiments.
4.3.1 Training language proficiency
To test the impact of training language proficiency, we select German and Chinese as representatives of high-resource languages, and Arabic and Hebrew as representatives of low-resource languages as training languages. Then, we train models on the KFRD in these languages and evaluated their performance across the 10 languages. The accuracy results are shown in Figure 5, and the transfer ratio is presented in Figure 5.
The results show that, though training with low-resource languages resulted in slightly lower accuracy (which is particularly noticeable in the arithmetic dataset), the models show no significant differences in transfer ratio when trained with high-resource or low-resource languages, indicating that the proficiency and resource of the training language has no significant effect on the cross-lingual transfer of knowledge-free reasoning.
4.3.2 Target language proficiency in specific directions
In previous experiments, we observe strong cross-linguistic transfer performance between most languages. However, the transferability from English to Arabic and Hebrew was significantly weaker. We hypothesize that this is related to the model’s weaker language proficiency in these two target languages.
In this section, we select models from Hugging Face that have undergone Continuous Pre-Training (CPT), Supervised Fine-Tuning (SFT), and a combination of both (CPT + SFT) on the LLaMA-2 or Mistral platforms. These adapted models have better proficiency in the respective languages. The selected models are listed in Table 2 in Appendix F.
We train a Vanilla model and the above fine-tuned models on the KFRD in English and test their transferability from English to the respective target languages. The transfer rate results are shown in Figure 6, and the accuracy results are provided in Figure 14 in Appendix G.
The experimental results indicate that the Vanilla model exhibits very low transfer rates for low-resource languages. However, after applying CPT, SFT, or CPT+SFT, the transfer rates increase significantly. Notably, for Hebrew, the transfer rate reached over 95%, approaching FCLT. This suggests that the proficiency in Arabic and Hebrew limits the cross-linguistic transfer of knowledge-free reasoning abilities from English to these languages. Improving proficiency in the target language can alleviate this limitation. This demonstrates that the cross-linguistic transfer of knowledge-free reasoning in specific directions is influenced by target language proficiency.
4.4 Analyzing differences in transfer characteristics
4.4.1 Overall computational similarity
In this section, we assess the LLaMA-2-7B-Chat model’s hidden states cosine similarity and Feed-Forward Network neuron activation overlap on knowledge-related tasks and knowledge-free reasoning tasks. We also evaluate changes in these metrics after fine-tuning with LoRA to understand the mechanisms supporting cross-lingual transfer of knowledge-free reasoning abilities. The experimental results are shown in Figures 7 and 8.
Figure 6: Averaged XLTR from English to Arabic/Hebrew on KFRD for models in different stages trained in Arabic/Hebrew
Figure 7: CS for different datasets in the LLaMA-2-7B-Chat model. The black lines on each bar indicate the 99% confidence intervals estimated with bootstrap sampling (Efron, 1992).
Figure 8: NAO for different dataset in the LLaMA-2-7B-Chat at activation thresholds ranging from 0.1 to 0.9. Shaded areas: 99% confidence intervals estimated with bootstrap sampling; Solid lines: results of the original model; Dashed lines: results of the LoRA tuned model. The meanings of the shaded areas and dashed lines in Figures 9 and 10 are consistent with those described here.
Internal representation of knowledge-free reasoning task is better aligned than retrieval
The results in Figure 7 indicate that the CS of the model on knowledge-free reasoning tasks is significantly higher than that on knowledge-related reasoning tasks both before and after training. Additionally, after training on the knowledge-free reasoning dataset, the model’s CS shows a significant increase. In contrast, after training on the knowledge-related reasoning dataset, the increase in CS is not significant and may even decrease. This suggests that adapting to knowledge-free reasoning tasks results in a more aligned hidden space processing across different languages, whereas adapting to knowledge-related reasoning tasks leads to relatively divergent hidden space processing.
Neuron activation pattern of knowledge-free reasoning task is more similar than retrieval
Neuron analysis further elucidates this phenomenon. The results in Figure 8 show that, across all threshold settings, NAO for knowledge-free reasoning tasks is significantly higher than for knowledge-related reasoning tasks. This indicates that the model tends to use similar neurons for processing knowledge-free reasoning tasks across different languages, resulting in similar neuron activation patterns and computational processes. Consistent with the hidden states results, after training on the knowledge-free reasoning dataset, the model’s NAO shows a significant increase, whereas there is no significant improvement and even a decline after training on the knowledge-related reasoning dataset. This suggests that training on knowledge-free reasoning tasks makes neuron activation characteristics across different languages more similar, whereas training on knowledge-related reasoning tasks makes these characteristics relatively divergent.
These results provide a comprehensive analysis of the differing cross-lingual transfer effectiveness between knowledge-free reasoning and knowledge-related reasoning from a computational similarity perspective. We hypothesize that this difference is because the model tends to store knowledge for different languages in different neurons, while using similar neuron groups for knowledge-free reasoning functions.
4.4.2 Layer-wise computational similarity
To analyze the relationship between hidden states and neuron activation characteristics with cross-lingual transfer of reasoning abilities more granularly, we performed a layer-wise analysis of the cross-lingual cosine similarity of hidden states and cross-lingual neural activation overlap. The experimental results are shown in Figures 9 and 10.
Figure 9: CS for different layers of the LLaMA-2-7B-Chat.
Figure 10: NAO for different layers of the LLaMA-2-7B-Chat at an activation threshold of 0.4.
It is observed that the phenomena identified in the previous section, specifically the significantly higher CS and NAO for knowledge-free reasoning tasks compared to knowledge tasks with more pronounced improvements after fine-tuning, are most evident in the middle layers (layers 6-25) of the model and less in the shallow layers (layers 1-6). Previous work (Zhao et al., 2024; Wendler et al., 2024) suggested that the middle layers of LLMs are primarily responsible for conceptual reasoning, which is cross-lingual. This hypothesis aligns with our findings and further supports the view that reasoning capabilities can transfer across languages.
Additionally, in the upper layers (layers 25-32) of the model, the CS and NAO for knowledge-free reasoning tasks are similar to those for knowledge tasks before training, but show significant improvement after training. This phenomenon is accompanied by a consistent performance improvement across different languages. The layered functionality theory of LLMs suggests that the upper layers are mainly responsible for generating the tokens of the target language responses. Therefore, we infer that task fine-tuning makes the model’s behavior more consistent during the decoding phase, resulting in better multilingual task performance.
5 Related Work
Multilingual reasoning evaluation
Laskar et al. (2023) performed evaluation for multilingual ability of ChatGPT. Shi et al. (2022) found LLMs can perform reasoning in multiple languages using CoT, even for those languages with very low resources.
Cross-lingual transfer
Gao et al. (2024) evaluated the cross-lingual transferability of models on multiple reasoning datasets, finding significant variations in transfer performance across different datasets. Additionally, Zhu et al. (2024) discovered that training on translated questions can enhance the cross-lingual transferability of reasoning tasks. Ye et al. (2023) assessed the imbalance of knowledge across different languages in LLMs, observing weak cross-lingual transferability of knowledge. Furthermore, Hu et al. (2024) found that knowledge transferability remains weak across various settings.
Analysis of multilingual internal representation
Zhao et al. (2024) analyzed the way LLMs handle multilingualism and suggested a three-phase working pattern, which includes understanding, task solving and generation. Wendler et al. (2024) also arrived at a similar conclusion.
6 Conclusion
In this study, we analyze the reasons behind the differing cross-lingual transfer abilities of LLMs on various reasoning tasks. We divide reasoning tasks into two components: knowledge retrieval and knowledge-free reasoning. We adapt several existing reasoning datasets and construct a knowledge-free reasoning dataset for analysis.
Experiments on these datasets demonstrated that the requirement of knowledge retrieval significantly lowers the cross-lingual transfer performance of reasoning tasks, while the knowledge-free reasoning ability can nearly fully transferred between languages.
In addition, we analyze the difference of computational processes between knowledge-related and knowledge-free reasoning tasks to explain the gap in cross-lingual transfer. The results show that the overall cross-lingual computational similarity of knowledge-free reasoning is significantly higher than that of knowledge-related reasoning, particularly in the middle-high layers of the model. This suggests that knowledge-free reasoning benefits from a shared neural mechanism across languages, whereas knowledge storage is more language-specific.
Limitations
One key limitation of this paper is the model selection and language coverage. In our exploration of language proficiency and interpretability experiments, we primarily rely on the LLaMA-2 model. Additionally, other parts of our research utilize only a few models, which may oversimplify the descriptions of model performance and behavior. In terms of language coverage, although we included ten languages from different language families, this number is still insufficient compared to the thousands of languages globally. This limitation is partly due to our computational resource constraints. With adequate resources, the proposed methods could be extended to other models and languages to further validate our conclusions.
Another limitation of our study is the depth of the interpretability analysis. We aim to investigate whether different knowledge-free reasoning tasks utilize the same neurons and whether knowledge is stored in different neurons for different languages. However, our support for this hypothesis is primarily based on macro-level numerical analyses, without precisely identifying specific reasoning neurons and knowledge neurons. This limitation restricts our fine-grained understanding of the model’s internal mechanisms. Future research should conduct more detailed neuron-level analyses to verify these hypotheses.
Ethics Statement
The authors declare no competing interests. All datasets used in this study are sourced from publicly available repositories and do not contain sensitive information, such as personal data. The data generated by GPT-4 have been verified to be non-toxic and are used exclusively for research purposes. The use of LLaMA-2 models, as well as several other large language models, complies with their respective licenses.
Appendix A Detailed Description of Knowledge-Free Reasoning
The KFRD is generated using a unified template, consisting entirely of multiple-choice questions with four options. Each question is structured into three parts: input, output, and transformation rules. Specific examples can be seen in Table 1, and the templates used for these examples are shown in Figure 11.
Arithmetic Reasoning
Input 11,645 (one or two numbers)
Transformation Rule Addition (a mathematical operation)
Output Options A) 595 B) 536 C) 771 D) 656 (one number, sometimes the result is two numbers)
Symbolic Reasoning
Input education, game, president, night, man (3-5 words in the corresponding language)
Transformation Rule Swap the positions of the 5th and 2nd words; Delete the 2nd word (symbolic operations within three layers)
Output Options A) education, president, night, game B) education, problem, night, game C) hand, president, night, game D) education, house, night, game
Logical Reasoning
Input Alex is Aurora Vale. Everything that is Aurora Vale is Omicron Delta. Stella is not Chronos Wasteland. Max is not Dreamweaver’s Haven. Suppose Sally is Whispering Meadows, then Sally is Chimerical Citadel. Everything that is Ebonwyrm Abyss is Phoenixfire Ridge. (8 propositions)
Transformation Rule Implication Elimination (a logical rule)
Output Options A) Alex is Seraphim Heights. B) Alex is Tempestwilds. C) Alex is Omicron Delta. D) Polly is Arcadia Reach.
Table 1: Examples of different tasks in the KFRD dataset
A.1 Arithmetic Reasoning
This dataset transforms two input numbers through mathematical operations into one or two output numbers. The mathematical operations include basic arithmetic operations (addition, subtraction, multiplication, division), equality, geometric progression, arithmetic progression, and sorting.
  • Input: Numbers are randomly generated within the range of 0-999.
  • Transformation rules: Each rule generates an equal number of samples.
  • Output: Generated through transformation rules, constrained within the range of 0-999. Other options are randomly generated, ensuring a single correct answer.
A.2 Symbolic Reasoning
This dataset transforms 3-5 input words from the corresponding language through symbolic operations to generate the output. Symbolic operations include repetition, addition, deletion, reordering, and their combinations. Considering that single-step symbolic operations are too simple, we chose up to three-step symbolic operations.
  • Input: Randomly select 3-5 words from a specific language. We chose 100 simple English words and translated them into other languages using Google Translate.
  • Transformation rules: The dataset includes equal amounts of single-step, two-step, and three-step symbolic operations. For single-step operations, each rule generates an equal number of samples. For two-step and three-step operations, rule combinations are randomly selected.
  • Output: Generated through transformation rules. Other options are partially randomly generated and partially based on random replacements from the original input, ensuring consistent length and a unique correct answer.
A.3 Logical Reasoning
This dataset generates output from a subset of eight input propositions using logical rules. Logical rules include Implication Elimination, Conjunction Introduction, Conjunction Elimination, Disjunction Introduction, Disjunction Elimination, and Proof by Contradiction. The Logical rules are referenced from Saparov et al. (2024).
  • Input: Eight propositions are generated using proposition templates and randomly selected entities, proposition templates referenced from Saparov et al. (2024) and entities from Saparov et al. (2024) and Gao et al. (2024). Missing languages were supplemented using Google Translate.
  • Transformation rules: Each logical rule generates an equal number of samples.
  • Output: Generated through logical rules. Other options are partially based on entities appearing in the propositions and partially randomly generated, ensuring a unique correct answer.
Instruction: The output is the result of applying a specific transformation rule to the input. In this question, you will be given an input value and its corresponding transformation rule. Based on this information, determine the correct output from the options provided: A, B, C, or D. Please give the corresponding answer option directly. Transformation Rule: {Transformation Rule} Input: {Input} Based on the above rule and input, choose the correct output from the following options: A. Output: {Output1} B. Output: {Output2} C. Output: {Output3} D. Output: {Output4} Answer:
Figure 11: Example prompt template for our KFRD dataset
Training Arabic Hebrew
Vanilla LLaMA-2-7B-Chat Mistral-7B-Instruct-v0.1
SFT Llama-2-7b-chat-arabic-lora (Icebear-AI, 2024) -
CPT SambaLingo-Arabic-Base (Csaki et al., 2024) DictaLM-2.0 (DICTA, 2024)
CPT+SFT SambaLingo-Arabic-Chat DictaLM-2.0-Instruct
Table 2: Training models for Arabic and Hebrew
Appendix B Dataset Examples
We provide some examples of adapted datasets used in this paper in Table 3.
StrategyQA
Question Are more people today related to Genghis Khan than Julius Caesar?
Facts 1. Julius Caesar had three children. 2. Genghis Khan had sixteen children. 3. Modern geneticists have determined that out of every 200 men today has DNA       that can be traced to Genghis Khan.
Answer True
MKQA
Query Who sings "I Hear You Knocking But You Can’t Come In"?
Answers Dave Edmunds
BoolQ
Question Do Iran and Afghanistan speak the same language?
Answer True
AmbigQA
Question How often does spermatogenesis—the production of sperm—occur?
Answer 74 days
Table 3: Examples of adapted datasets used in this paper
Appendix C Language Choice
This section provides an overview of the languages utilized in our research, highlighting the primary countries where they are spoken and their respective language families. Refer to Table 4 for detailed information.
ISO Countries Language Family
en US, UK Germanic
de Germany, Austria Germanic
fr France, Canada Romance
it Italy Romance
pl Poland Slavic
ru Russia, Belarus Slavic
ar Egypt, Algeria Afro-Asiatic
he Israel Afro-Asiatic
ja Japan Japonic
zh China (Mainland) Sino-Tibetan
Table 4: Correspondence between Languages, Countries, and Language Families
Appendix D Implementation Details for Interpretability
D.1 Dataset adjustments
We made slight modifications of the datasets to ensure that the last character of different datasets is uniformly a "?" token. Since we are analyzing the internal representation of the last token, in this way, we can eliminate interference caused by the inconsistent input token, which may make the representation unreliable, especially in the bottom layers. Another reason why we append the token "?" is that it can act as a trigger to let the model start the process of preparing to answer the question, which is what we are analyzing.
For knowledge-free reasoning dataset, we added the phrase "Which option should I choose?" in different languages. For the MKQA and BoolQ datasets, where some questions did not end with a "?", we added a "?". All other datasets already ended with a "?".
D.2 Calculation method for activation values
We use the output of the gate linear layer in the SwiGLU module of the LLaMA model, processed through the SiLU function, as the activation values.
Dataset Samples Epoch
StrategyQA 2061 4
KFRD Arithmetic 8000 4
KFRD Symbolic 2000 1
KFRD Logical 4000 1
Table 5: Training epoch and number of samples of fine-tuned datasets in the transferability experiments
Dataset Samples
StrategyQA 228
KFRD Arithmetic 800
KFRD Symbolic 500
KFRD Logical 500
Table 6: The size of testset used in the transferability experiments
Appendix E Experiments Details
This section outlines the details of our experiments for reproducibility.
E.1 Infrastructure
We used the following scientific artifacts in our research:
  • PyTorch (Ansel et al., 2024, BSD license), a framework for building and running deep learning models.
  • Transformers (Wolf et al., 2020, Apache-2.0 license), a library providing a user friendly interface for running and fine-tuning pre-trained models.
  • DeepSpeed (Rasley et al., 2020, Apache-2.0 license), a library optimizing the parallel training of the deep learning models.
  • LLaMA-Factory (Zheng et al., 2024, Apache-2.0 license), a library that provides a unifying way to easily fine-tune large language models with parameter efficient fine-tuning technique like LoRA.
E.2 Hyperparameters
In the fine-tuning of all models, we use a learning rate of 2e-4 with a cosine learning rate scheduler. We clip the gradient norm to 1.0, use a total batch size of 64, set the rank of LoRA to 128, and alpha to 16. The LoRA adapters are applied to all the linear layers within Transformer blocks.
The numbers of training epoch and samples used in the transferability experiments are listed in Table 5. These numbers are tuned to enable LLaMA-2-7B-Chat to achieve 85% + accuracy on the corresponding tasks. The size of testsets used in the transferability experiments are shown in Table 6.
In the interpretability experiments, we adjust the number of training epochs or the size of the syntactic datasets to keep the number of total update steps on all datasets around 150, which avoids interference of different update steps on experimental results. We report the average cosine similarity and neuron activation overlap of 100 samples from each dataset.
E.3 Computation resources
All the fine-tuning experiments can be done on 4 NVIDIA Tesla V100 32GB GPUs. Each fine-tuning can be done in no more than 2 hours.
Appendix F Models Used in the Target Language Proficiency Experiment
The continue pre-training or fine-tuning models of LLaMA-2-7B and Mistral-7B used in the target language proficiency experiment in 4.3.2 are listed in Table 2.
Appendix G Additional Results of Experiment
Here we provide the accuracy of the above experiments in Figure 12, 13 and 14.
Figure 12: Left: Accuracy of different models on StrategyQA. Solid and dashed line represent the result of With Facts and No Facts setting, respectively. Middle: Accuracy of different models on StrategyQA before fine-tuning. Right: Accuracy of LLaMA-2-7B-Chat on StrategyQA under various settings. The translucent line represents the accuracy before finetuning on the specific tasks (which are all around 50%).
Figure 13: Accuracy of various models on different parts of KFRD. The translucent line represents the accuracy before finetuning on the specific tasks.
Figure 14: Averaged Accuracy on English and Arabic/Hebrew KFRD for models in different stages trained in Arabic/Hebrew

### v2

@@
 However, the relationship between capabilities in different languages is less explored.
-In this work, we decompose the process of reasoning tasks into two separated parts: knowledge retrieval and knowledge-free reasoning, and analyze the cross-lingual transferability of them.
-With adapted and constructed knowledge-free reasoning datasets, we show that the knowledge-free reasoning capability can be nearly perfectly transferred across various source-target language directions despite the secondary impact of resource in some specific target languages, while cross-lingual knowledge retrieval significantly hinders the transfer.
-Moreover, by analyzing the hidden states and feed-forward network neuron activation during the reasoning tasks, we show that higher similarity of hidden representations and larger overlap of activated neurons could explain the better cross-lingual transferability of knowledge-free reasoning than knowledge retrieval.
-Thus, we hypothesize that knowledge-free reasoning embeds in some language-shared mechanism, while knowledge is stored separately in different languages.
+In this work, we decompose the process of reasoning tasks into two separated components: knowledge retrieval and knowledge-free reasoning, and analyze the relationship between cross-lingual transferability and these two components.
+With adapted commonsense reasoning datasets and constructed knowledge-free reasoning datasets, we show that the knowledge-free reasoning capability can be nearly perfectly transferred across various source-target language directions despite the secondary impact of resource in some specific target languages, while cross-lingual knowledge retrieval significantly hinders the transfer.
+Moreover, by analyzing the hidden states and feed-forward network neuron activation during the reasoning, we show that higher similarity of hidden representations and larger overlap of activated neurons could explain the better cross-lingual transferability of knowledge-free reasoning than knowledge retrieval.
+Thus, we hypothesize that knowledge-free reasoning shares similar neurons in different languages for reasoning, while knowledge is stored separately in different languages.
  [^1] Our code and data is available at: https://github.com/NJUNLP/Knowledge-Free-Reasoning.
@@
 1 Introduction
-Large language models (LLMs) today have shown strong multitask and multilingual performance in various domains Huang and Chang (2022), including robust reasoning capabilities across multiple languages Ranaldi and Zanzotto (2023), even for low-resource languages in the training corpus Shi et al.
+Large language models (LLMs) today have shown strong multitask and multilingual performance in various domains Huang and Chang (2022), including robust reasoning capabilities across multiple languages Ranaldi et al.
+(2024), even for low-resource languages in the training corpus Shi et al.
 (2022).
-Previous study reveals that these multilingual LLMs possess certain ability of multilingual transfer Ye et al.
+Previous study reveals that these multilingual LLMs possess certain ability of multilingual transfer Qi et al.
+(2023); Gao et al.
+(2024); Ye et al.
 (2023), which means the skills or knowledge learned with one language can be automatically transferred to another language without extra training.
 However, the effect of such cross-lingual transfer varies across tasks.
-In knowledge-related tasks, current LLMs show unsatisfactory cross-lingual transfer Gao et al.
-(2024), while in certain reasoning tasks, more effective transfer is observed Qi et al.
+In certain tasks, especially knowledge retrieval tasks, current LLMs show unsatisfactory cross-lingual transfer Qi et al.
+(2023); Gao et al.
+(2024), while in certain reasoning tasks, more effective transfer is observed Ye et al.
 (2023).
 Previous study lacks the analysis on the difference between these tasks, and does not dig further into the specific factors affecting the transfer effectiveness.
-In this study, we divide a reasoning task into two separated components: knowledge retrieval and knowledge-free reasoning, and investigate the transfer effectiveness of them in current LLMs.
-The former means recalling information encountered during training, while the latter refers to organizing the given knowledge in the context into a final answer.
-Based on such definition, we perform evaluation and interpretability analysis on LLaMA-2 models Touvron et al.
-(2023) to compare the cross-lingual transferability of knowledge retrieval and knowledge-free reasoning.
-In the evaluation part, by controlling the requirement of knowledge retrieval in the StrategyQA (Geva et al., 2021) dataset and constructing a knowledge-free reasoning dataset, we compare the cross-lingual transfer effectiveness of knowledge-related and knowledge-free reasoning.
-In the interpretability analysis part, we assess the cross-lingual computational similarity of hidden states and Feed-Forward Network (FFN) neural activation to trace the computational process of cross-lingual transfer of knowledge-related and knowledge-free reasoning in LLMs.
+Figure 1: Cross-lingual transfer involves training a model in one language and evaluating it in another.
+In this figure, the scenario depicts training in English.
+Reasoning tasks encompass both knowledge retrieval and knowledge-free reasoning.
+The cross-lingual transfer ratio is significantly lower for knowledge retrieval tasks (e.g., ZH case in EN: "Crocodiles, alligators, and pigeons are dangerous animals") compared to knowledge-free reasoning tasks, which transfer well across languages (e.g., ZH case in EN: "22 plus 23 equals 45").
+In this study, we divide a general reasoning task into two separated components: knowledge retrieval and knowledge-free reasoning.
+The former means recalling certain information from pre-training, while the latter refers to organizing the given knowledge in the context to perform inference and derive a final answer.
+Figure 1 provides a clearer understanding of these two components and illustrates the cross-lingual transfer issues explored in this paper.
+This paper includes both an evaluation part and an interpretability analysis part.
+In the evaluation part, we focus on the impact of knowledge retrieval component on cross-lingual transfer in reasoning tasks, and the transferability of knowledge-free reasoning capability, by adapting existing reasoning datasets with different levels of knowledge retrieval demand and creating a clean knowledge-free reasoning dataset, which only includes the knowledge-free reasoning component.
+In the interpretability analysis part, we assess the cross-lingual computational similarity of hidden states and Feed-Forward Network neuron activation to trace and compare the computational process of cross-lingual transfer of knowledge retrieval and knowledge-free reasoning components in LLMs.
 Our main findings are:
-  • Knowledge retrieval significantly hinders cross-lingual transfer. The more knowledge retrieval is required in the task, the lower effectiveness of cross-lingual transfer is observed.
-  • The knowledge-free reasoning ability can be effectively transferred to other languages after finetuning in one, while the model’s ability in these languages are less important.
-  • The overall cross-lingual computational similarity of knowledge-free reasoning is significantly higher than knowledge-related reasoning.
-Also, the difference in such similarity level is the most significant in the middle-high layers of the model.
-This suggests a language-shared mechanism of knowledge-free reasoning in multilingual LLMs.
-2 Methods
-The work in this paper consists of two parts: evaluation and interpretability analysis. For each part, we employ specific datasets and design certain metrics or measurements.
+  • Retrieval component significantly hinders cross-lingual transfer of reasoning tasks.
+The more knowledge retrieval is required in the task, the lower effectiveness of cross-lingual transfer is observed.
+  • The ability of knowledge-free reasoning component can be near-perfectly transferred to other languages after fine-tuning in one, while the model’s language proficiency in the target languages is also important.
+  • The overall cross-lingual computational similarity for knowledge-free reasoning tasks is significantly higher than for knowledge retrieval tasks, especially in the middle-high layers, which are primarily used for reasoning (Zhao et al., 2024; Wendler et al., 2024).
+This suggests a language-shared reasoning mechanism in multilingual LLMs.
+2 Evaluation Methodology
+2.1 Overview
+Our evaluation focuses on two main aspects:
+Impact of Knowledge Retrieval Demand on Cross-Lingual Transfer in Reasoning Tasks
+We aim to analyze how varying levels of knowledge retrieval demand affect cross-lingual transfer in reasoning tasks.
+For this purpose, we leverage the commonsense reasoning datasets that provide questions along with several facts required to answer them.
+By controlling the number of facts provided to the model within the context, we can manipulate the levels of demand for knowledge retrieval.
+As more facts are provided, the model relies less on its internal knowledge storage.
+This controlled setup enables us to analyze how the demand for knowledge retrieval influences the cross-lingual transfer of the overall reasoning abilities.
+Cross-Lingual Transfer of Knowledge-Free Reasoning
+We also aim to investigate the cross-lingual transfer of knowledge-free reasoning, which is less explored in previous work.
+However, existing reasoning datasets often contain some degree of knowledge retrieval.
+For instance, while StrategyQA provides knowledge required for reasoning, it is not always complete.
+Similarly, certain mathematical datasets, like ASDiv, require knowledge retrieval for some problems (as demonstrated in Appendix G).
+This introduces noise when evaluating the cross-lingual transfer of knowledge-free reasoning.
+To address this, we constructed a new dataset, the Knowledge-Free Reasoning Dataset (KFRD), which entirely eliminates the need for knowledge retrieval.
+In addition, we selected several existing datasets that, to the best extent possible, meet the requirements of knowledge-free reasoning to further validate our conclusions.
+A more detailed explanation of why we constructed KFRD and the dataset selection criteria can be found in Appendix G.
 2.2 Datasets
-This part introduces the datasets used in this paper for the analysis of the cross-lingual transfer of reasoning tasks.
-For datasets available only in English or part of the tested languages, we translate them into other languages with Google Translate.
-2.1.1 Adapted StrategyQA dataset
-The StrategyQA dataset is a multi-domain reasoning dataset including common sense, arithmetic, symbolic, and logical reasoning, where model is given at least two pieces of evidence from Wikipedia, and required to give "yes" or "no" answers to the questions.
-An example can be found in Table 3 in Appendix B.
-Since Wikipedia is a commonly used source in LLM pretraining, the model can either collect the required information from evidence in the context, or retrieve it from its internal knowledge strorage.
-Thus, by masking out some of the evidence, we can control the need of knowledge retrieval in the reasoning process, which allows us to analyze the impact of knowledge retrieval on the cross-lingual transfer of reasoning skills.
-Namely, we design two kinds of scenarios in the experiments:
-  • No Fact (NF): The model is given only the questions and must retrieve the necessary knowledge (evidence).
+This section introduces the datasets used for evaluation. More details on the datasets and the construction process are in Appendix A.
+2.2.1 Reasoning dataset with variable knowledge retrieval demand
+We adapt the popular commonsense reasoning datasets, StrategyQA (Geva et al., 2021) and QASC (Khot et al., 2020), to analyze the impact of knowledge retrieval on cross-lingual transfer.
+They provide pieces of evidence from Wikipedia for answering the question.
+Examples can be found in Table A5.
+Namely, we design two kinds of scenarios with variable knowledge retrieval demand in the experiments:
+  • No Fact (NF): The model is given only the questions.
   • With Fact (WF): The model is provided with the questions and some of the evidence.
@@
 2.2.2 Knowledge-free reasoning dataset
-Although the need for knowledge retrieval can be controlled with the adapted StrategyQA dataset, the questions themselves may still contain implicit requirement for knowledge retrieval.
-To rule out such effect and focus on the knowledge-free reasoning process, we construct a data that minimizes the need for knowledge retrieval.
 Inspired by Wei et al.
-(2022)’s taxonomy of reasoning tasks, we developed the Knowledge-Free Reasoning Dataset (KFRD), which consists of three fundamental reasoning tasks: arithmetic reasoning, symbolic reasoning, and logic reasoning in the 4-choice format.
-The dataset is constructed by filling GPT-4 Achiam et al.
-(2023) generated entities or simple English words into a unified template with one input part, one transformation rule, and one options part.
-Then, using the translated templates, entities and words, the dataset is expanded to all the tested languages.
-The template and examples can be found in Figure 11 and Table 1 in Appendix A.
-Arithmetic Reasoning:
-transforming one or two input numbers through mathematical operations into one or two output numbers.
-The mathematical operations include: basic arithmetic operations (addition, subtraction, multiplication, division), equality, geometric progression, arithmetic progression, and sorting.
-Symbolic Reasoning:
-transforming 3-5 words from the respective language input through symbolic operations to obtain the output.
-The symbolic operations include: repetition, addition, deletion, reordering, and their combinations, with up to three operations applied in combination.
-Logical Reasoning:
-selecting propositions derived from eight given premises using provided logical rules.
-The logical rules include: Implication Elimination, Conjunction Introduction, Conjunction Elimination, Disjunction Introduction, Disjunction Elimination, and Proof by Contradiction.
-2.1.3 Knowledge-related reasoning dataset
-To further compare the cross-lingual transfer of knowledge retrieval and knowledge-free reasoning, three knowledge-related reasoning datasets are used in the interpretability analysis.
-The three datasets are:
-MKQA:
-The MKQA dataset Longpre et al.
-(2021) is an open-domain question-answering evaluation set designed to provide a challenging benchmark for multilingual question-answering.
-The dataset consists of 10,000 question-answer pairs and covers 26 different languages.
-BoolQ:
-BoolQ (Clark et al., 2019) is a question answering dataset for yes/no questions, comprising 15,942 examples.
-The questions originate from queries to the Google search engine with Wikipedia pages as answers.
-AmbigQA:
-AmbigQA (Min et al., 2020) is an open-domain dataset for question answering that includes 14,042 questions. It addresses the issue of ambiguous answers by rephrasing the related questions.
-Some examples of these datasets are shown in Table 3 in Appendix B.
+(2022)’s taxonomy of reasoning tasks, we developed the KFRD, which consists of three fundamental reasoning tasks: arithmetic reasoning (e.g., addition, subtraction, and other mathematical operations), symbolic reasoning(e.g., deletion, reordering, and other symbolic operations), and logic reasoning(e.g., Implication Elimination and other basic logical rules) .
+It is designed to evaluate a broad spectrum of knowledge-free reasoning and cross-lingual transfer performance.
+We utilized GPT-4 Achiam et al.
+(2023) to generate multilingual parallel templates and fictitious entities, followed by manual verification.
+We then used code to generate the dataset based on these templates, entities, and predefined rules.
+This approach ensures that the tasks can be completed without requiring additional knowledge and guarantees the correctness of the QA pairs.
+The templates are multiple-choice questions, each composed of one input part, one transformation rule, and one options part.
+The examples and template are provided in Table 1 and Figure A1.
+We also use the ASDiv Miao et al.
+(2021), Coin Flip Wei et al.
+(2022), and ProofWriter Tafjord et al.
+(2020) dataset as a representation of arithmetic, symbolic, and logical reasoning to further validate our conclusions.
+Arithmetic Reasoning
+Input 11, 645 (two numbers)
+Transformation Rule Addition (a mathematical operation)
+Output Options A) 595  B) 536 C) 771  D) 656
+Symbolic Reasoning
+Input education, game, president, night, man (3-5 words in the corresponding language)
+Transformation Rule Swap the positions of the 5th and 2nd words; Delete the 2nd word (1-3 symbolic operations)
+Output Options A) education, president, night, game  B) education, problem, night, game C) hand, president, night, game  D) education, house, night, game
+Logical Reasoning
+Input Alex is Aurora Vale.
+Everything that is Aurora Vale is Omicron Delta.
+Stella is not Chronos Wasteland.
+Max is not Dreamweaver’s Haven.
+Suppose Sally is Whispering Meadows, then Sally is Chimerical Citadel.
+Everything that is Ebonwyrm Abyss is Phoenixfire Ridge.
+(6 propositions)
+Transformation Rule Implication Elimination (a logical rule)
+Output Options A) Alex is Seraphim Heights.  B) Alex is Tempestwilds. C) Alex is Omicron Delta.  D) Polly is Arcadia Reach.
+Table 1: Examples of different tasks in the KFRD dataset
 2.3 Evaluation metric
-In order to measure the model’s cross-lingual transferability while mitigating the difficulty difference in of the datasets, we calculate the Cross-lingual Transfer Ratio (XLTR) following the prior work Gao et al.
-(2024).
-The formula is as follows:
+In order to assess the model’s cross-lingual transferability, we select the Cross-lingual Transfer Ratio (XLTR) as the evaluation metric, following Gao et al. (2024). The formula is as follows:
 $\text{XLTR}(s,t)=(\frac{|C_{s}\cap C_{t}|}{|C_{s}|}-A_{r})/(1-A_{r})$
@@
 $C_{x}$ represents the set of correct answers in language $x$, and $A_{r}$ is the accuracy of random choices for the given task.
-If after training on a task in one specific language (namely, English), the model shows an XLTR score close to 100% in other untrained languages, we say it achieves Fully Cross-Lingual Transfer (FCLT) in this task.
+If the model shows an XLTR score close to 100% in a language direction, we say it achieves fully cross-lingual transfer in this direction.
+We also evaluate the accuracy of models before fine-tuning on all datasets and find poor performance, suggesting that most of the model’s ability on transferred languages stem from cross-lingual transfer.
+3 Experiment Settings
+3.1 Language and model choice
+Language choice
+To capture linguistic diversity, we selected ten languages based on their extensive use and representation of diverse linguistic families, following Gao et al.
+(2024).
+The languages selected are English (en), German (de), French (fr), Italian (it), Russian (ru), Polish (pl), Arabic (ar), Hebrew (he), Chinese (zh), and Japanese (ja).
+Unless specified, we fine-tune the model in English and evaluate it on other languages.
+Further details are provided in Appendix B.
+Model choice
+We selected several LLMs, including LLaMA-2-7B-Chat (Touvron et al., 2023), BLOOMZ-MT-7B (Muennighoff et al., 2023), Mistral-7B-Instruct-v0.1 (Jiang et al., 2023), and Qwen-1.5-7B-Chat (Bai et al., 2023), for our experiments.
+To optimize resource use and demonstrate results clearly, we used LLaMA-2-7B-Chat (Touvron et al., 2023) as a representative model for some analyses.
+3.2 Fine-tuning and decoding settings
+We perform LoRA fine-tuning (Hu et al., 2021) on all model blocks in all experiments due to the limited computational resources. More details about fine-tuning can be found in Appendix D.
+For decoding, we use constrained decoding in all experiments to ensure the model generates only the desired options (e.g., Yes/No for StrategyQA, A/B/C/D for KFRD).
+4 Results
+Figure 2: XLTR of different models on StrategyQA. Solid lines: WF-all results; Dashed lines: NF results. The label of training language (en) is capitalized.
+Figure 3: XLTR of LLaMA-2-7B-Chat on StrategyQA under different settings.
+Figure 4: XLTR on the different parts of KFRD
+Figure 5: XLTR on the existing pseudo knowledge-free reasoning datasets
+Figure 6: XLTR of LLaMA-2-7B-Chat on three parts of KFRD. The different lines indicate different trained languages.
+4.1 Impact of knowledge retrieval demand on cross-lingual transfer
+We analyze the impact of the amount of knowledge retrieve on cross-lingual transfer in different settings of the reasoning dataset.
+The results of StrategyQA for the cross-lingual transfer ratio are presented in Figure 2, while the accuracy results are detailed in Figure A2.
+Knowledge retrieval requirement harms cross-lingual transfer
+The experimental results indicate that, for all languages, the cross-lingual transfer ratio of all models are significantly higher when the necessary knowledge for reasoning is provided compared to when it is not.
+This suggests that the requirement for knowledge retrieval significantly hinders the model’s cross-lingual transferability when solving reasoning tasks.
+More knowledge retrieval lowers cross-lingual transfer
+We further conduct detailed evaluations using the LLaMA-2-7B-Chat model to observe the changes in cross-lingual transfer ratios as the amount of knowledge provided varies.
+As shown in Figure 3, the experimental results demonstrate that the transfer ratio decreases as the demand for knowledge retrieval increases.
+This further validates the conclusion that the retrieval of more knowledge significantly impacts cross-lingual transferability.
+The results on the QASC dataset were consistent with those mentioned above. Detailed results can be found in Figure A3 and A4.
+4.2 The cross-lingual transfer of knowledge-free reasoning
+We assess the cross-lingual transferability of the model’s knowledge-free reasoning capabilities by evaluating the performance on KFRD and three corresponding existing reasoning datasets.
+The resulting cross-lingual transfer ratios are shown in Figures 4 and 5, while the accuracy results are presented in Figures A5 and A5.
+The results demonstrate that the KFRD exhibits extremely high cross-lingual transfer performance for most language pairs.
+For 7 out of the 9 languages, it can be observed that the cross-lingual transfer ratio in knowledge-free reasoning tasks often exceeds 90%, with some instances approaching 100%, thus achieving near-full cross-lingual transfer.
+Moreover, results from three existing datasets further validate this finding, showing that all models achieve satisfactory transfer ratios across high-resource languages.
+For some low-resource languages, such as Hebrew and Arabic in LLaMA-2, German and Hebrew in BLOOMZ  [^2] For the language distribution, please see Appendix F., the cross-lingual transferability is significantly poorer.
+We hypothesize that this may be due to the model’s weaker language proficiency in these languages, which negatively impacts its transferability.
+Further analysis of this issue is provided in the following section.
+It is noticeable that there are still minor differences in XLTR between KFRD and the existing datasets in the arithmetic reasoning and logical reasoning tasks.
+However, these differences do not affect the overall conclusion.
+We manually check the data samples and find that there are some interfering cases that can affect cross-lingual transfer, such as minor knowledge retrieves, translation issues, and counterfactual information, as detailed discussed in the Appendix G.
+4.3 Impact of language proficiency on cross-lingual transfer
+4.3.1 Training language proficiency
+To evaluate the impact of training language proficiency, based on the data distribution of LLaMA-2 (see Appendix F) and previous experiments, we selected German and Chinese as representatives of high-resource languages, and Arabic and Hebrew as representatives of low-resource languages for training.
+Then, we train models on the KFRD in these languages and evaluated their performance across the 10 languages.
+The resulting cross-lingual transfer ratios are presented in Figure 6, while the accuracy results are shown in Figure A6.
+The results show that the models show no significant differences in transfer ratio when trained with high-resource or low-resource languages, indicating that the proficiency and resource of the training language has no significant effect on the cross-lingual transfer of knowledge-free reasoning.
+4.3.2 Target language proficiency
+In previous experiments, we observe the transferability from English to Arabic and Hebrew was significantly weaker in LLaMA-2 and Mistral.
+We hypothesize that this is related to the model’s weaker language proficiency in these two target languages.
+In this section, we select models from Hugging Face that have undergone Continual Pre-Training (CPT), Supervised Fine-Tuning (SFT), and a combination of both (CPT + SFT) on the LLaMA-2 or Mistral platforms.
+These adapted models have better proficiency in the respective languages.
+The selected models are listed in Table A1.
+The transfer ratio results of the vanilla and the above fine-tune models are shown in Figure 7, and the accuracy results are provided in Figure A7.
+Figure 7: Averaged XLTR from English to Arabic/Hebrew across three parts of our KFRD dataset for models in different stages trained in Arabic/Hebrew
+The results indicate that the vanilla model exhibits very low transfer rates for low-resource languages.
+However, after applying CPT, SFT, or CPT+SFT, the transfer ratio increases significantly.
+Notably, for Hebrew, the transfer ratio reach over 95%, achieving fully cross-lingual transfer.
+This suggests that proficiency in Arabic and Hebrew limits the cross-linguistic transfer of the knowledge-free reasoning component, while improving proficiency in the target language can alleviate this limitation.
+5 Interpretability Analysis
+5.1 Overview
+Built on previous research (Hu et al., 2024; Gao et al., 2024) and our experiments, we observed that the cross-lingual transferability of knowledge retrieval ability is significantly weaker than that of knowledge-free reasoning.
+To better understand the reasons behind this difference, we conducted a detailed analysis on model internals using two methods: Cosine Similarity of Hidden States and Neuron Activation.
+Both of the methods have been widely used to measure text similarity (Reimers and Gurevych, 2019; Malkiel et al., 2022; Wang et al., 2024) and analyze models (Dalvi et al., 2019; Sajjad et al., 2022; Rai and Yao, 2024).
 5.2 Interpretability measurements
-We employ two metrics to evaluate the cross-lingual computational consistency of the LLM during knowledge-related and knowledge-free reasoning processes.
+This section introduces the measurements used for interpretability analysis. Further details for these metrics are in Appendix C.
 5.2.1 Cosine similarity of hidden states (CS)
@@
 5.2.2 Neuron Activation Overlap (NAO)
-The Neuron Activation Overlap refers to the extent of overlap between activated neurons across multiple languages for the same corpus. The calculation steps are as follow:
-  1. 1.
-    Input a question in multiple languages into the model and extract the activation values of the last token in the question in each language.
-  2. 2.
-    Identify activated neurons. A neuron is considered activated if the absolute value of its activation exceeds a certain threshold.
-  3. 3.
-    Calculate the overlap of activated neurons in all the tested languages.
-For a question sample $x$, the formula is as follows:
+Neuron Activation Overlap measures the extent of shared neuron activations across languages for the same input.
+To calculate NAO, we input a question in multiple languages, extract the neuron activation values of the last token of the input, and identify the neurons whose absolute values surpass a set threshold, labeling them as activated.
+Then their overlap (NAO) is computed as follows for a question sample $x$:
 $\text{NAO}(x)=\frac{\left|\mathcal{L}\right|\cdot\left|\bigcap_{l\in\mathcal{L}}S^{l}(x)\right|}{\sum_{l\in\mathcal{L}}\left|S^{l}(x)\right|}$
 where $\mathcal{L}$ is set of languages, and $S^{l}(x)$ is the set of activated neurons on sample $x$ in language $l$. After that, the NAO of all tested samples are averaged to report the final score.
-3 Experiment Settings
-3.1 Language and model choice
-Language choice
-To capture a broad spectrum of linguistic diversity, we selected ten languages based on their extensive use and representation of diverse linguistic families, following Gao et al.
-(2024).
-The languages selected are English (en), German (de), French (fr), Italian (it), Russian (ru), Polish (pl), Arabic (ar), Hebrew (he), Chinese (zh), and Japanese (ja).
-Further details are provided in Appendix C.
-Model choice
-We selected several LLMs including LLaMA-2-7B-Chat (Touvron et al., 2023), BLOOMZ-MT-7B (Muennighoff et al., 2023), Mistral-7B-Instruct-v0.1 (Jiang et al., 2023), and Qwen-1.5-7B-Chat (Bai et al., 2023) to conduct transferability experiments on existing and synthetic datasets.
-For experiments focusing on the impact of the training language (Section 4.3.1) and interpretability (Section 4.4), we used LLaMA-2-7B-Chat as a representative model for analysis.
-For the experiments evaluating impact of the target language proficiency in specific directions (Section 4.3.2), we selected derived models of LLaMA-2-7B and Mistral-7B to conduct experiments on Arabic and Hebrew respectively, both of which are low-resource languages.
-3.2 Fine-tuning settings
-We perform LoRA fine-tuning (Hu et al., 2021) on all model blocks in all experiments due to the limited computational resources. More details about fine-tuning can be found in Appendix E.
-3.3 Decoding settings
-In all experiments, we perform constrained decoding to prevent the model from generating tokens other than the desired choices (Yes/No for StrategyQA, A/B/C/D for KFRD).
-In the interpretability analysis, we use the last token of the question to collect the hidden states and neural activation values, because the last input token is used to predict the next token, it gradually incorporates the primary information of the entire sentence, reflecting the overall thought process for the entire problem (Meng et al., 2022; Stolfo et al., 2023; Wu et al., 2024).
-By focusing on the model’s computational pathway for reasoning rather than calculating the similarity between multilingual sentences, we can better understand how the model processes complex queries.
-Calculating with an output token, on the other hand, would make it difficult to interpret the reasoning process.
-Additionally, token counts differ across languages, complicating direct comparisons.
-Therefore, using the last input token helps in standardizing the analysis across different languages.
-To ensure consistency in the final token across different datasets, we added a language-specific ’?’ to certain datasets. For details on these modifications and metrics, please refer to Appendix D.
-4 Results
-Figure 1: XLTR of different models on StrategyQA. Solid lines: WF results; Dashed lines: NF results. The label of training language (en) is capitalized.
-Figure 2: XLTR of LLaMA-2-7B-Chat on StrategyQA under different settings.
-NF stands for the No Facts setting, while WF-1 and WF-2 indicate that one or two facts are given for each question, respectively.
-WF-all signifies that all facts are provided.
-Figure 3: XLTR on the different parts of KFRD
-Figure 4: Accuracy of LLaMA-2-7B-Chat from different languages on three parts of KFRD
-Figure 5: XLTR of LLaMA-2-7B-Chat from different languages on three parts of KFRD
-4.1 Impact of knowledge retrieval on cross-lingual transfer
-We analyze the impact of the amount of retrieved knowledge on cross-lingual transfer in different settings of the StrategyQA dataset.
-The results for cross-lingual transfer ratio are presented in Figure 1, while the accuracy results are detailed in Figure 12 in Appendix G.
-Knowledge retrieval requirement harms cross-lingual transfer
-The experimental results indicate that, for all languages, the cross-lingual transfer ratio of all models are significantly higher when the necessary knowledge for reasoning is provided compared to when it is not.
-This suggests that the requirement for knowledge retrieval significantly hinders the model’s cross-lingual transferability when solving reasoning problems.
-Less knowledge retrieval makes better cross-lingual transfer
-Additionally, we conduct a more detailed evaluation using the LLaMA-2-7B-Chat model to observe the changes in cross-lingual transfer ratio with varying amounts of knowledge entries.
-As shown in Figure 2, the experimental results demonstrate that the transfer rate increases as the number of provided knowledge entries increases from none to all required knowledge.
-This further validates the conclusion that the retrieval of more knowledge significantly impacts cross-lingual transferability.
-4.2 The cross-lingual transfer of knowledge-free reasoning
-We assess the cross-lingual transferability of the model’s knowledge-free reasoning capabilities.
-We evaluate the performance of the KFRD by training in English and transferring to the other 9 languages.
-The resulting cross-lingual transfer ratio are shown in Figure 3, and accuracy results are shown in Figure 13 in Appendix G.
-The results demonstrate that the KFRD exhibits extremely high cross-lingual transfer performance for most language pairs.
-For 7 out of the 9 languages, it can be observed that the cross-lingual transfer performance in knowledge-free reasoning tasks often exceeds 90%, with some instances approaching 100%, thereby achieving near FCLT.
-However, for the two lowest-resource languages Arabic and Hebrew, the cross-lingual transferability is significantly poorer.
-We hypothesize that this may be due to the model’s weaker language proficiency in these two languages, which adversely affects its transferability.
-Further analysis on this issue is provided in the next section.
-4.3 Impact of language proficiency on cross-lingual transfer
-In this section, we explore the impact of language proficiency of the training and target language with additional experiments.
-4.3.1 Training language proficiency
-To test the impact of training language proficiency, we select German and Chinese as representatives of high-resource languages, and Arabic and Hebrew as representatives of low-resource languages as training languages.
-Then, we train models on the KFRD in these languages and evaluated their performance across the 10 languages.
-The accuracy results are shown in Figure 5, and the transfer ratio is presented in Figure 5.
-The results show that, though training with low-resource languages resulted in slightly lower accuracy (which is particularly noticeable in the arithmetic dataset), the models show no significant differences in transfer ratio when trained with high-resource or low-resource languages, indicating that the proficiency and resource of the training language has no significant effect on the cross-lingual transfer of knowledge-free reasoning.
-4.3.2 Target language proficiency in specific directions
-In previous experiments, we observe strong cross-linguistic transfer performance between most languages.
-However, the transferability from English to Arabic and Hebrew was significantly weaker.
-We hypothesize that this is related to the model’s weaker language proficiency in these two target languages.
-In this section, we select models from Hugging Face that have undergone Continuous Pre-Training (CPT), Supervised Fine-Tuning (SFT), and a combination of both (CPT + SFT) on the LLaMA-2 or Mistral platforms.
-These adapted models have better proficiency in the respective languages.
-The selected models are listed in Table 2 in Appendix F.
-We train a Vanilla model and the above fine-tuned models on the KFRD in English and test their transferability from English to the respective target languages.
-The transfer rate results are shown in Figure 6, and the accuracy results are provided in Figure 14 in Appendix G.
-The experimental results indicate that the Vanilla model exhibits very low transfer rates for low-resource languages.
-However, after applying CPT, SFT, or CPT+SFT, the transfer rates increase significantly.
-Notably, for Hebrew, the transfer rate reached over 95%, approaching FCLT.
-This suggests that the proficiency in Arabic and Hebrew limits the cross-linguistic transfer of knowledge-free reasoning abilities from English to these languages.
-Improving proficiency in the target language can alleviate this limitation.
-This demonstrates that the cross-linguistic transfer of knowledge-free reasoning in specific directions is influenced by target language proficiency.
-4.4 Analyzing differences in transfer characteristics
+5.3 Knowledge retrieval dataset
+We selected MKQA Longpre et al.
+(2021), BoolQ (Clark et al., 2019), and AmbigQA (Min et al., 2020) as representative datasets of knowledge retrieval tasks for the interpretability analysis.
+Most questions in these datasets can be answered through a single instance of knowledge retrieval.
+Examples of these datasets are shown in Table A8.
+5.4 Interpretability results
 5.4.1 Overall computational similarity
-In this section, we assess the LLaMA-2-7B-Chat model’s hidden states cosine similarity and Feed-Forward Network neuron activation overlap on knowledge-related tasks and knowledge-free reasoning tasks.
-We also evaluate changes in these metrics after fine-tuning with LoRA to understand the mechanisms supporting cross-lingual transfer of knowledge-free reasoning abilities.
-The experimental results are shown in Figures 7 and 8.
-Figure 6: Averaged XLTR from English to Arabic/Hebrew on KFRD for models in different stages trained in Arabic/Hebrew
-Figure 7: CS for different datasets in the LLaMA-2-7B-Chat model. The black lines on each bar indicate the 99% confidence intervals estimated with bootstrap sampling (Efron, 1992).
+In this section, we assess the original and fine-tuned LLaMA-2-7B-Chat model’s CS and NAO on knowledge retrieval and knowledge-free reasoning tasks.
+The experimental results are shown in Figures 8 and 9.
+Figure 8: CS for different datasets in the LLaMA-2-7B-Chat model. Black lines on each bar indicate the 99% confidence intervals estimated with bootstrap sampling (Efron, 1992).
 Figure 9: NAO for different dataset in the LLaMA-2-7B-Chat at activation thresholds ranging from 0.1 to 0.9.
 Shaded areas: 99% confidence intervals estimated with bootstrap sampling; Solid lines: results of the original model; Dashed lines: results of the LoRA tuned model.
-The meanings of the shaded areas and dashed lines in Figures 9 and 10 are consistent with those described here.
-Internal representation of knowledge-free reasoning task is better aligned than retrieval
-The results in Figure 7 indicate that the CS of the model on knowledge-free reasoning tasks is significantly higher than that on knowledge-related reasoning tasks both before and after training.
-Additionally, after training on the knowledge-free reasoning dataset, the model’s CS shows a significant increase.
-In contrast, after training on the knowledge-related reasoning dataset, the increase in CS is not significant and may even decrease.
-This suggests that adapting to knowledge-free reasoning tasks results in a more aligned hidden space processing across different languages, whereas adapting to knowledge-related reasoning tasks leads to relatively divergent hidden space processing.
-Neuron activation pattern of knowledge-free reasoning task is more similar than retrieval
+The meanings of the shaded areas and dashed lines in Figures 10 and 11 are consistent with those described here.
+Internal representation of knowledge-free reasoning task is better aligned than knowledge retrieval
+The results in Figure 8 indicate that the CS of the model on knowledge-free reasoning tasks is significantly higher than that on knowledge retrieval tasks both before and after fine-tuning.
+Additionally, after fine-tuning on knowledge-free reasoning datasets, the CS increases significantly on the corresponding datasets, while fine-tuning on knowledge retrieval datasets shows no significant improvement and may even lead to a decrease.
+This suggests that adapting to knowledge-free reasoning tasks results in a more aligned hidden space processing across languages.
+Neuron activation pattern of knowledge-free reasoning task is more similar than knowledge retrieval
 Neuron analysis further elucidates this phenomenon.
-The results in Figure 8 show that, across all threshold settings, NAO for knowledge-free reasoning tasks is significantly higher than for knowledge-related reasoning tasks.
-This indicates that the model tends to use similar neurons for processing knowledge-free reasoning tasks across different languages, resulting in similar neuron activation patterns and computational processes.
-Consistent with the hidden states results, after training on the knowledge-free reasoning dataset, the model’s NAO shows a significant increase, whereas there is no significant improvement and even a decline after training on the knowledge-related reasoning dataset.
-This suggests that training on knowledge-free reasoning tasks makes neuron activation characteristics across different languages more similar, whereas training on knowledge-related reasoning tasks makes these characteristics relatively divergent.
-These results provide a comprehensive analysis of the differing cross-lingual transfer effectiveness between knowledge-free reasoning and knowledge-related reasoning from a computational similarity perspective.
-We hypothesize that this difference is because the model tends to store knowledge for different languages in different neurons, while using similar neuron groups for knowledge-free reasoning functions.
+The results in Figure 9 show that, across all activation threshold settings, NAO for knowledge-free reasoning tasks is significantly higher than for knowledge retrieval tasks.
+This indicates that the model tends to use similar neurons for processing knowledge-free reasoning tasks across different languages, resulting in similar neuron activation patterns.
+Consistent with the hidden states results, after training on the knowledge-free reasoning dataset, NAO increases significantly, whereas there is no significant improvement and even a decline after training on the knowledge retrieval dataset.
+This suggests that training on knowledge-free reasoning tasks makes neuron activation characteristics across different languages more similar, leading to the conclusion that the knowledge-free reasoning ability share a similar set of neurons.
+These results provide a comprehensive analysis of the different cross-lingual transfer effectiveness between knowledge-free reasoning and knowledge retrieval component from a computational similarity perspective.
+We hypothesize that this difference is because the model stores knowledge for different languages in different neurons, while using similar neuron groups for knowledge-free reasoning.
 5.4.2 Layer-wise computational similarity
-To analyze the relationship between hidden states and neuron activation characteristics with cross-lingual transfer of reasoning abilities more granularly, we performed a layer-wise analysis of the cross-lingual cosine similarity of hidden states and cross-lingual neural activation overlap.
-The experimental results are shown in Figures 9 and 10.
+To gain finer-grained insights, we performed a layer-wise analysis of CS and NAO. The experimental results are shown in Figures 10 and 11.
 Figure 10: CS for different layers of the LLaMA-2-7B-Chat.
 Figure 11: NAO for different layers of the LLaMA-2-7B-Chat at an activation threshold of 0.4.
-It is observed that the phenomena identified in the previous section, specifically the significantly higher CS and NAO for knowledge-free reasoning tasks compared to knowledge tasks with more pronounced improvements after fine-tuning, are most evident in the middle layers (layers 6-25) of the model and less in the shallow layers (layers 1-6).
+It is observed that the significantly higher CS and NAO for knowledge-free reasoning tasks, compared to knowledge retrieval tasks, are most pronounced in the middle layers (layers 6-25).
 Previous work (Zhao et al., 2024; Wendler et al., 2024) suggested that the middle layers of LLMs are primarily responsible for conceptual reasoning, which is cross-lingual.
-This hypothesis aligns with our findings and further supports the view that reasoning capabilities can transfer across languages.
-Additionally, in the upper layers (layers 25-32) of the model, the CS and NAO for knowledge-free reasoning tasks are similar to those for knowledge tasks before training, but show significant improvement after training.
-This phenomenon is accompanied by a consistent performance improvement across different languages.
-The layered functionality theory of LLMs suggests that the upper layers are mainly responsible for generating the tokens of the target language responses.
-Therefore, we infer that task fine-tuning makes the model’s behavior more consistent during the decoding phase, resulting in better multilingual task performance.
+This hypothesis aligns with our findings and further supports the view that knowledge-free reasoning capabilities can transfer across languages.
+Additionally, the upper layers (26-32) show similar CS and NAO patterns for both knowledge-free reasoning and knowledge retrieval tasks before training, but training improvements are only notable in knowledge-free reasoning.
+We find that fine-tuning on knowledge-free tasks significantly enhances multilingual accuracy, leading to more consistent outputs.
+Since the upper layers primarily handle token generation (Zhao et al., 2024; Wendler et al., 2024), this consistency improvement results in higher CS and NAO.
 6 Related Work
@@
 (2022) found LLMs can perform reasoning in multiple languages using CoT, even for those languages with very low resources.
+Their analysis mainly evaluated different reasoning tasks, but did not investigate the reasons for performance variations.
 Cross-lingual transfer
@@
 (2024) evaluated the cross-lingual transferability of models on multiple reasoning datasets, finding significant variations in transfer performance across different datasets.
-Additionally, Zhu et al.
-(2024) discovered that training on translated questions can enhance the cross-lingual transferability of reasoning tasks.
-Ye et al.
-(2023) assessed the imbalance of knowledge across different languages in LLMs, observing weak cross-lingual transferability of knowledge.
 Furthermore, Hu et al.
 (2024) found that knowledge transferability remains weak across various settings.
+Building on their conclusions, we distinguish between the knowledge retrieval and knowledge-free reasoning components and extend the analysis to all reasoning tasks.
+We include further discussion of related reasoning works in Appendix H.
 Analysis of multilingual internal representation
@@
 (2024) also arrived at a similar conclusion.
-6 Conclusion
+Expanding on their findings, we further analyzed the differences in how LLMs handle reasoning and knowledge tasks across languages.
+7 Conclusion and Discussion
 In this study, we analyze the reasons behind the differing cross-lingual transfer abilities of LLMs on various reasoning tasks.
 We divide reasoning tasks into two components: knowledge retrieval and knowledge-free reasoning.
-We adapt several existing reasoning datasets and construct a knowledge-free reasoning dataset for analysis.
-Experiments on these datasets demonstrated that the requirement of knowledge retrieval significantly lowers the cross-lingual transfer performance of reasoning tasks, while the knowledge-free reasoning ability can nearly fully transferred between languages.
-In addition, we analyze the difference of computational processes between knowledge-related and knowledge-free reasoning tasks to explain the gap in cross-lingual transfer.
-The results show that the overall cross-lingual computational similarity of knowledge-free reasoning is significantly higher than that of knowledge-related reasoning, particularly in the middle-high layers of the model.
-This suggests that knowledge-free reasoning benefits from a shared neural mechanism across languages, whereas knowledge storage is more language-specific.
+Our experiments demonstrated that the demand for knowledge retrieval significantly hinders the cross-lingual transfer performance, while the knowledge-free reasoning ability can be nearly fully transferred between languages.
+This discrepancy arises because knowledge-free reasoning relies on shared neural mechanisms across languages, while knowledge storage tends to be more language-specific.
+Based on these findings, for knowledge, we recommend prioritizing the inclusion of multilingual data in training corpora in the future.
+For reasoning, emphasis should be placed on the quality of reasoning data rather than the number of languages.
+Furthermore, for future multilingual analysis, we recommend investigating knowledge retrieval and knowledge-free reasoning components individually to gain more targeted and detailed insights.
 Limitations
@@
 The use of LLaMA-2 models, as well as several other large language models, complies with their respective licenses.
-Appendix A Detailed Description of Knowledge-Free Reasoning
-The KFRD is generated using a unified template, consisting entirely of multiple-choice questions with four options.
+Appendix A Detailed of Dataset
+A.1 Detailed description of Knowledge-Free Reasoning Dataset
+The KFRD is generated using a unified template, consisting entirely of multi-choice questions with four options.
+We first create parallel templates for 10 languages using GPT-4 and then fill in different parts of the template with pre-defined rules.
 Each question is structured into three parts: input, output, and transformation rules.
-Specific examples can be seen in Table 1, and the templates used for these examples are shown in Figure 11.
-Arithmetic Reasoning
-Input 11,645 (one or two numbers)
-Transformation Rule Addition (a mathematical operation)
-Output Options A) 595 B) 536 C) 771 D) 656 (one number, sometimes the result is two numbers)
-Symbolic Reasoning
-Input education, game, president, night, man (3-5 words in the corresponding language)
-Transformation Rule Swap the positions of the 5th and 2nd words; Delete the 2nd word (symbolic operations within three layers)
-Output Options A) education, president, night, game B) education, problem, night, game C) hand, president, night, game D) education, house, night, game
-Logical Reasoning
-Input Alex is Aurora Vale.
-Everything that is Aurora Vale is Omicron Delta.
-Stella is not Chronos Wasteland.
-Max is not Dreamweaver’s Haven.
-Suppose Sally is Whispering Meadows, then Sally is Chimerical Citadel.
-Everything that is Ebonwyrm Abyss is Phoenixfire Ridge.
-(8 propositions)
-Transformation Rule Implication Elimination (a logical rule)
-Output Options A) Alex is Seraphim Heights. B) Alex is Tempestwilds. C) Alex is Omicron Delta. D) Polly is Arcadia Reach.
-Table 1: Examples of different tasks in the KFRD dataset
-A.1 Arithmetic Reasoning
+Specific examples can be seen in Table 1, and the templates used for these examples are shown in Figure A1.
+A.1.1 Arithmetic reasoning
 This dataset transforms two input numbers through mathematical operations into one or two output numbers.
-The mathematical operations include basic arithmetic operations (addition, subtraction, multiplication, division), equality, geometric progression, arithmetic progression, and sorting.
+The mathematical operations include addition, subtraction, multiplication, division, equality, geometric progression, arithmetic progression, and sorting.
+Each of the three parts are generated by the following rules:
   • Input: Numbers are randomly generated within the range of 0-999.
@@
   • Output: Generated through transformation rules, constrained within the range of 0-999. Other options are randomly generated, ensuring a single correct answer.
-A.2 Symbolic Reasoning
+A.1.2 Symbolic reasoning
 This dataset transforms 3-5 input words from the corresponding language through symbolic operations to generate the output.
@@
 Considering that single-step symbolic operations are too simple, we chose up to three-step symbolic operations.
+Each of the three parts are generated by the following rules:
   • Input: Randomly select 3-5 words from a specific language. We chose 100 simple English words and translated them into other languages using Google Translate.
@@
 Other options are partially randomly generated and partially based on random replacements from the original input, ensuring consistent length and a unique correct answer.
-A.3 Logical Reasoning
+A.1.3 Logical reasoning
 This dataset generates output from a subset of eight input propositions using logical rules.
@@
 (2024).
+Each of the three parts are generated by the following rules:
   • Input: Eight propositions are generated using proposition templates and randomly selected entities, proposition templates referenced from Saparov et al.
@@
 Output: {Output4} Answer:
-Figure 11: Example prompt template for our KFRD dataset
+Figure A1: Example prompt template for our KFRD dataset
 Training Arabic Hebrew
@@
 CPT+SFT SambaLingo-Arabic-Chat DictaLM-2.0-Instruct
-Table 2: Training models for Arabic and Hebrew
-Appendix B Dataset Examples
-We provide some examples of adapted datasets used in this paper in Table 3.
-StrategyQA
-Question Are more people today related to Genghis Khan than Julius Caesar?
-Facts 1. Julius Caesar had three children. 2. Genghis Khan had sixteen children. 3. Modern geneticists have determined that out of every 200 men today has DNA       that can be traced to Genghis Khan.
-Answer True
-MKQA
-Query Who sings "I Hear You Knocking But You Can’t Come In"?
-Answers Dave Edmunds
-BoolQ
-Question Do Iran and Afghanistan speak the same language?
-Answer True
-AmbigQA
-Question How often does spermatogenesis—the production of sperm—occur?
-Answer 74 days
-Table 3: Examples of adapted datasets used in this paper
-Appendix C Language Choice
+Table A1: Training models for Arabic and Hebrew
+A.2 Detail of existing pseudo knowledge-free reasoning datasets
+Here we provide more details on the datasets used in the experiment.
+A.3 Translation process for English-only datasets
+For datasets available only in English, we translate them into other languages with Google Translate and verify translation quality with GPT-4.
+Google Translate is highly regarded in the field of commercial translation and is widely used in multilingual research (Chen et al., 2021; Ye et al., 2023; Omar and Salih, 2024; Song et al., 2024).
+To ensure translation accuracy, we sampled a subset of translation results and employed GPT-4 for verification.
+Due to budget constraints, we were unable to employ human translators.
+For the StrategyQA dataset, we utilized Google Translate and conducted a sample check of 100 items for each language using GPT-4.
+This process resulted in an overall quality score of 4.47 (on a scale of 1-5), which we consider acceptable for our purposes.
+Appendix B Language Choice
 This section provides an overview of the languages utilized in our research, highlighting the primary countries where they are spoken and their respective language families.
-Refer to Table 4 for detailed information.
-ISO Countries Language Family
+Refer to Table A2 for detailed information.
+ISO Country Samples Language Family
 en US, UK Germanic
@@
 zh China (Mainland) Sino-Tibetan
-Table 4: Correspondence between Languages, Countries, and Language Families
-Appendix D Implementation Details for Interpretability
-D.1 Dataset adjustments
-We made slight modifications of the datasets to ensure that the last character of different datasets is uniformly a "?" token.
+Table A2: Correspondence between Languages, Country Samples, and Language Families
+Appendix C Implementation Details for Interpretability
+C.1 Calculation method for activation values
+We use the output of the gate linear layer in the SwiGLU module of the LLaMA model, processed through the SiLU function, as the activation values.
+C.2 Reasons for using the last token for analysis
+In the interpretability analysis, we use the last token of the question to collect the hidden states and neural activation values, because the last input token is used to predict the next token, it gradually incorporates the primary information of the entire sentence, reflecting the overall thought process for the entire problem (Meng et al., 2022; Stolfo et al., 2023; Wu et al., 2024).
+By focusing on the model’s computational pathway for reasoning rather than calculating the similarity between multilingual sentences, we can better understand how the model processes complex queries.
+Calculating with an output token, on the other hand, would make it difficult to interpret the reasoning process.
+Additionally, token counts differ across languages, complicating direct comparisons.
+Therefore, using the last input token helps in standardizing the analysis across different languages.
+C.3 Dataset adjustments
+To ensure consistency in the final token across different datasets, we made slight modifications by adding a language-specific "?" where needed.
 Since we are analyzing the internal representation of the last token, in this way, we can eliminate interference caused by the inconsistent input token, which may make the representation unreliable, especially in the bottom layers.
@@
 All other datasets already ended with a "?".
-D.2 Calculation method for activation values
-We use the output of the gate linear layer in the SwiGLU module of the LLaMA model, processed through the SiLU function, as the activation values.
 Dataset Samples Epoch
@@
 KFRD Logical 4000 1
-Table 5: Training epoch and number of samples of fine-tuned datasets in the transferability experiments
+Table A3: Training epoch and number of samples of fine-tuned datasets in the transferability experiments
 Dataset Samples
@@
 KFRD Logical 500
-Table 6: The size of testset used in the transferability experiments
-Appendix E Experiments Details
+Table A4: The size of testset used in the transferability experiments
+Appendix D Experiments Details
 This section outlines the details of our experiments for reproducibility.
-E.1 Infrastructure
+D.1 Infrastructure
 We used the following scientific artifacts in our research:
@@
   • LLaMA-Factory (Zheng et al., 2024, Apache-2.0 license), a library that provides a unifying way to easily fine-tune large language models with parameter efficient fine-tuning technique like LoRA.
-E.2 Hyperparameters
+D.2 Hyperparameters
 In the fine-tuning of all models, we use a learning rate of 2e-4 with a cosine learning rate scheduler.
@@
 The LoRA adapters are applied to all the linear layers within Transformer blocks.
-The numbers of training epoch and samples used in the transferability experiments are listed in Table 5.
+The numbers of training epoch and samples used in the transferability experiments are listed in Table A3.
 These numbers are tuned to enable LLaMA-2-7B-Chat to achieve 85% + accuracy on the corresponding tasks.
-The size of testsets used in the transferability experiments are shown in Table 6.
+The size of testsets used in the transferability experiments are shown in Table A4.
 In the interpretability experiments, we adjust the number of training epochs or the size of the syntactic datasets to keep the number of total update steps on all datasets around 150, which avoids interference of different update steps on experimental results.
 We report the average cosine similarity and neuron activation overlap of 100 samples from each data set.
-E.3 Computation resources
+StrategyQA
+Question Are more people today related to Genghis Khan than Julius Caesar?
+Facts 1. Julius Caesar had three children. 2. Genghis Khan had sixteen children. 3. Modern geneticists have determined that out of every 200 men today has DNA       that can be traced to Genghis Khan.
+Answer Yes
+QASC
+Question Climate is generally described in terms of what?
+Facts 1. Climate is generally described in terms of temperature and moisture. 2. Fire behavior is driven by local weather conditions such as winds, temperature and       moisture.
+Options A. sand B. occurs over a wide range C. forests D. Global warming E. rapid changes occur F. local weather conditions G. measure of motion H. city life
+Answer F
+Table A5: Examples of knowledge-involved datasets
+D.3 Computation resources
 All the fine-tuning experiments can be done on 4 NVIDIA Tesla V100 32GB GPUs. Each fine-tuning can be done in no more than 2 hours.
-Appendix F Models Used in the Target Language Proficiency Experiment
-The continue pre-training or fine-tuning models of LLaMA-2-7B and Mistral-7B used in the target language proficiency experiment in 4.3.2 are listed in Table 2.
-Appendix G Additional Results of Experiment
-Here we provide the accuracy of the above experiments in Figure 12, 13 and 14.
-Figure 12: Left: Accuracy of different models on StrategyQA.
+D.4 Models used in the target language proficiency experiment
+The continue pre-training or fine-tuning models of LLaMA-2-7B and Mistral-7B used in the target language proficiency experiment in 4.3.2 are listed in Table A1.
+Appendix E Additional Results of Experiment
+Here we provide the accuracy of the above experiments in Figure A2, A5, A6 and A7.
+We provide the results of the QASC dataset in Figure A3 and A4. The results show that the more knowledge provided leads to better cross-lingual transferability, which aligns with our conclusion.
+Appendix F Language Distribution of Model Training Corpora
+In this section, we present the language distribution of the pre-training corpora, referencing Table A6 from the LLaMA2 paper (Touvron et al., 2023) and Table A7 from the BLOOM paper (Workshop et al., 2022).
+Unfortunately, we were unable to locate the corresponding distribution data for Mistral and Qwen.
+For LLaMA2, languages such as Arabic and Hebrew were not included in the table, indicating that their proportions are lower than 0.005%, categorizing them as extremely low-resource languages.
+The other eight languages discussed in the paper are represented.
+Notably, German and Chinese rank as high-resource languages, accounting for 0.17% and 0.13% of the corpus, respectively, holding the second and fifth highest positions.
+For BLOOM, only English, French, Chinese, and Arabic are explicitly listed, while other languages are not reported in the table.
+Language Percent Language Percent
+en 89.70% uk 0.07%
+unknown 8.38% ko 0.06%
+de 0.17% ca 0.04%
+fr 0.16% sr 0.04%
+sv 0.15% id 0.03%
+zh 0.13% cs 0.03%
+es 0.13% fi 0.03%
+ru 0.13% hu 0.03%
+nl 0.12% no 0.03%
+it 0.11% ro 0.03%
+ja 0.10% bg 0.02%
+pl 0.09% da 0.02%
+pt 0.09% sl 0.01%
+vi 0.08% hr 0.01%
+Table A6: Language distribution in pretraining data with percentage >= 0.005%.
+Most data is in English, meaning that Llama 2 will perform best for English-language use cases.
+The large unknown category is partially made up of programming code data.
+Language Size in Bytes Language Size in Bytes
+Akan 70,1554 Arabic 74,854,900,600
+Assamese 291,522,098 Bambara 391,747
+Basque 2,360,470,848 Bengali 18,606,823,104
+Catalan 17,792,493,289 Chichewa 1,187,405
+chiShona 6,638,639 Chitumbuka 170,360
+English 484,953,009,124 Fon 2,478,546
+French 208,242,620,434 Gujarati 1,199,986,460
+Hindi 24,622,119,985 Igbo 14,078,521
+Indonesian 19,972,325,222 isiXhosa 14,304,074
+isiZulu 8,511,561 Kannada 2,098,453,560
+Kikuyu 359,615 Kinyarwanda 40,428,299
+Kirundi 3,272,550 Lingala 1,650,804
+Luganda 4,568,367 Malayalam 3,662,571,498
+Marathi 1,775,483,122 Nepali 2,551,307,393
+Northern Sotho 1,764,506 Odia 1,157,100,133
+Portuguese 79,277,543,375 Punjabi 1,572,109,752
+Sesotho 751,034 Setswana 1,502,200
+Simplified Chinese 261,019,433,892 Spanish 175,098,365,045
+Swahili 236,482,543 Tamil 7,989,206,220
+Telugu 29,934,071,59 Traditional Chinese 762,489,150
+Twi 1,265,041 Urdu 2,781,329,959
+Vietnamese 43,709,279,959 Wolof 3,606,973
+Xitsonga 707,634 Yoruba 89,695,835
+Programming Languages 174,700,245,772
+Table A7: Linguistic makeup of the ROOTS corpus. The ROOTS corpus is a multilingual dataset used for pretraining the BLOOM model.
+Appendix G Reasons for Creating a New Dataset
+The primary reason for creating a new dataset is that most existing datasets involve knowledge retrieval, which does not align with our focus on knowledge-free reasoning.
+For instance, in StrategyQA, while necessary reasoning knowledge is provided, it may be incomplete.
+StrategyQA Example:
+  • Question: Are you likely to find a crucifix in Karachi?
+  • Facts: The crucifix is a symbol of Christianity. The vast majority of Pakistan’s population is Muslim.
+  • Missing Knowledge: It is not specified that Karachi is in Pakistan.
+Similarly, most existing math datasets also require knowledge retrieval to answer questions, such as the ASDiv-a dataset.
+ASDiv-a Example 1:
+  • Question: At the school’s book fair, Sam bought 13 adventure books and 17 mystery books. If 15 of the books were used, how many new books did he buy?
+  • Missing Knowledge: The new books are those that were not used.
+ASDiv-a Example 2:
+  • EN-Question: After the aviary was the zoo’s swamp area. Penny counted a total of 55 tree frogs, 10 poison frogs, and 13 wood frogs. How many frogs was Penny able to count?
+  • FR-Question: Après la volière se trouvait la zone marécageuse du zoo.
+Penny a dénombré un total de 55 rainettes, 10 grenouilles venimeuses et 13 grenouilles des bois.
+Combien de grenouilles Penny était-elle capable de compter ?
+  • Missing Knowledge: In English, it can be inferred that “poison frogs," “wood frogs," and “tree frogs" are all “frogs." However, in French, it is not directly inferable that “rainettes" are a type of “grenouilles," requiring additional knowledge retrieval.
+Some existing logic datasets are not designed with knowledge-free reasoning in mind, as they use real-world entities.
+This leads to situations where, although it is theoretically possible to answer without retrieving external knowledge, the retrieval of such knowledge might influence the final answer.
+For example, given the statement “Harry is a cat," the model might infer “Harry is an animal" based on its existing knowledge, without requiring contextual reasoning rules.
+Similarly, based on prior knowledge, the model might incorrectly assume “The squirrel likes the squirrel" as related, especially when the actual context is irrelevant.
+This issue becomes more pronounced when translation is involved.
+For instance, when translating from English to Chinese, “The squirrel likes the squirrel" may become “squirrels like squirrels," as Chinese does not use articles.
+This can amplify the influence of pre-existing knowledge, leading to incorrect answers.
+By constructing our own dataset, we also avoid potential translation issues that arise when existing datasets are used in different languages, ensuring that reasoning tasks are uniformly understood across languages.
+Another advantage of creating a new dataset is that we can control the difficulty level.
+If the dataset is too difficult and models have low accuracy in English, it would be meaningless to measure cross-lingual transferability.
+Moreover, a new dataset allows for a more comprehensive coverage of reasoning operations.
+Figure A2: Left: Accuracy of different models on StrategyQA.
 Solid and dashed line represent the result of With Facts and No Facts setting, respectively.
@@
 The translucent line represents the accuracy before finetuning on the specific tasks (which are all around 50%).
-Figure 13: Accuracy of various models on different parts of KFRD. The translucent line represents the accuracy before finetuning on the specific tasks.
-Figure 14: Averaged Accuracy on English and Arabic/Hebrew KFRD for models in different stages trained in Arabic/Hebrew
+Figure A3: XLTR of different models on QASC. Solid lines: WF-2 results; Dashed lines: NF results.
+Figure A4: XLTR of LLaMA-2-7B-Chat on QASC. Here WF-2 equals to WF-all, as QASC only has two pieces of evidence per sample.
+Figure A5: Accuracy of various models on different parts of KFRD. The translucent line represents the accuracy before finetuning on the specific tasks.
+Figure A6: Accuracy of LLaMA-2-7B-Chat on three parts of KFRD. The different lines indicate different trained languages.
+Figure A7: Averaged Accuracy on English and Arabic/Hebrew KFRD for models in different stages trained in Arabic/Hebrew
+MKQA
+Query Who sings "I Hear You Knocking But You Can’t Come In"?
+Answers Dave Edmunds
+BoolQ
+Question Do Iran and Afghanistan speak the same language?
+Answer True
+AmbigQA
+Question How often does spermatogenesis—the production of sperm—occur?
+Answer 74 days
+Table A8: Examples of adapted datasets used in this paper
+Appendix H Additional Works on Cross-lingual Transfer
+Devlin (2018) introduced mBERT, advancing cross-lingual transfer by capturing shared linguistic patterns in a unified embedding space, enabling zero-shot transfer without parallel corpora.
+Similarly, Conneau (2019) showed XLM’s effectiveness in optimizing multilingual embeddings, improving performance in translation and classification tasks.
+Ansell et al.
+(2021) proposed composable sparse fine-tuning, selectively fine-tuning sparse parameters across languages to reduce interference and boost performance, especially in low-resource settings, outperforming adapter-based methods in tasks like NER and NLI.
+Additionally, Ye et al.
+(2023) assessed the imbalance of knowledge across different languages in LLMs, observing weak cross-lingual transferability of knowledge.
+Zhu et al.
+(2024) discovered that training on translated questions can enhance the cross-lingual transferability of reasoning tasks.

### v3

@@
 Thus, we hypothesize that knowledge-free reasoning shares similar neurons in different languages for reasoning, while knowledge is stored separately in different languages.
- [^1] Our code and data is available at: https://github.com/NJUNLP/Knowledge-Free-Reasoning.
 Large Language Models Are Cross-Lingual Knowledge-Free Reasoners
@@
 The cross-lingual transfer ratio is significantly lower for knowledge retrieval tasks (e.g., ZH case in EN: "Crocodiles, alligators, and pigeons are dangerous animals") compared to knowledge-free reasoning tasks, which transfer well across languages (e.g., ZH case in EN: "22 plus 23 equals 45").
+Intuitively, abstract reasoning is considered a language-agnostic ability, and thus it should transfer well across languages.
+The observed performance differences in general reasoning tasks across languages may arise from other factors.
 In this study, we divide a general reasoning task into two separated components: knowledge retrieval and knowledge-free reasoning.
-The former means recalling certain information from pre-training, while the latter refers to organizing the given knowledge in the context to perform inference and derive a final answer.
+The former means recalling factual knowledge from pre-training  [^1] Similar to closed-book QA tasks, while the latter refers to organizing the given knowledge in the context to perform inference and derive a final answer [^2] Such as Coin Flip Wei et al.
+(2022).
 Figure 1 provides a clearer understanding of these two components and illustrates the cross-lingual transfer issues explored in this paper.
@@
 However, existing reasoning datasets often contain some degree of knowledge retrieval.
-For instance, while StrategyQA provides knowledge required for reasoning, it is not always complete.
-Similarly, certain mathematical datasets, like ASDiv, require knowledge retrieval for some problems (as demonstrated in Appendix G).
+For instance, while StrategyQA (Geva et al., 2021) provides knowledge required for reasoning, it is not always complete.
+Similarly, certain mathematical datasets, like ASDiv, require knowledge retrieval for some problems (as demonstrated in Appendix H).
 This introduces noise when evaluating the cross-lingual transfer of knowledge-free reasoning.
@@
 In addition, we selected several existing datasets that, to the best extent possible, meet the requirements of knowledge-free reasoning to further validate our conclusions.
-A more detailed explanation of why we constructed KFRD and the dataset selection criteria can be found in Appendix G.
+A more detailed explanation of why we constructed KFRD and the dataset selection criteria can be found in Appendix H.
 2.2 Datasets
@@
 2.2.1 Reasoning dataset with variable knowledge retrieval demand
-We adapt the popular commonsense reasoning datasets, StrategyQA (Geva et al., 2021) and QASC (Khot et al., 2020), to analyze the impact of knowledge retrieval on cross-lingual transfer.
+We adapt the popular commonsense reasoning datasets, StrategyQA and QASC (Khot et al., 2020), to analyze the impact of knowledge retrieval on cross-lingual transfer.
 They provide pieces of evidence from Wikipedia for answering the question.
@@
 4.1 Impact of knowledge retrieval demand on cross-lingual transfer
-We analyze the impact of the amount of knowledge retrieve on cross-lingual transfer in different settings of the reasoning dataset.
+We analyze the impact of the amount of knowledge retrieved on cross-lingual transfer in different settings of the reasoning dataset.
 The results of StrategyQA for the cross-lingual transfer ratio are presented in Figure 2, while the accuracy results are detailed in Figure A2.
@@
 Moreover, results from three existing datasets further validate this finding, showing that all models achieve satisfactory transfer ratios across high-resource languages.
-For some low-resource languages, such as Hebrew and Arabic in LLaMA-2, German and Hebrew in BLOOMZ  [^2] For the language distribution, please see Appendix F., the cross-lingual transferability is significantly poorer.
+For some low-resource languages, such as Hebrew and Arabic in LLaMA-2, German and Hebrew in BLOOMZ  [^3] For the language distribution, please see Appendix G., the cross-lingual transferability is significantly poorer.
 We hypothesize that this may be due to the model’s weaker language proficiency in these languages, which negatively impacts its transferability.
@@
 However, these differences do not affect the overall conclusion.
-We manually check the data samples and find that there are some interfering cases that can affect cross-lingual transfer, such as minor knowledge retrieves, translation issues, and counterfactual information, as detailed discussed in the Appendix G.
+We manually check the data samples and find that there are some interfering cases that can affect cross-lingual transfer, such as minor knowledge retrieves, translation issues, and counterfactual information, as detailed discussed in the Appendix H.
+We also evaluate the LLaMA-2-7B-Chat model on MMLU before and after finetuning, in order to address the concerns of over-fitting on the fine-tuned datasets and forgetting the world knowledge, which is detailed in Appendix F.
 4.3 Impact of language proficiency on cross-lingual transfer
 4.3.1 Training language proficiency
-To evaluate the impact of training language proficiency, based on the data distribution of LLaMA-2 (see Appendix F) and previous experiments, we selected German and Chinese as representatives of high-resource languages, and Arabic and Hebrew as representatives of low-resource languages for training.
+To evaluate the impact of training language proficiency, based on the data distribution of LLaMA-2 (see Appendix G) and previous experiments, we selected German and Chinese as representatives of high-resource languages, and Arabic and Hebrew as representatives of low-resource languages for training.
 Then, we train models on the KFRD in these languages and evaluated their performance across the 10 languages.
@@
 5.1 Overview
-Built on previous research (Hu et al., 2024; Gao et al., 2024) and our experiments, we observed that the cross-lingual transferability of knowledge retrieval ability is significantly weaker than that of knowledge-free reasoning.
+Built on previous research (Hu et al., 2024b; Gao et al., 2024) and our experiments, we observed that the cross-lingual transferability of knowledge retrieval ability is significantly weaker than that of knowledge-free reasoning.
 To better understand the reasons behind this difference, we conducted a detailed analysis on model internals using two methods: Cosine Similarity of Hidden States and Neuron Activation.
@@
 Furthermore, Hu et al.
-(2024) found that knowledge transferability remains weak across various settings.
+(2024a) found that knowledge transferability remains weak across various settings.
 Building on their conclusions, we distinguish between the knowledge retrieval and knowledge-free reasoning components and extend the analysis to all reasoning tasks.
-We include further discussion of related reasoning works in Appendix H.
+Additionally, Ye et al.
+(2023) assessed the imbalance of knowledge across different languages in LLMs, observing weak cross-lingual transferability of knowledge.
+Zhu et al.
+(2024) discovered that training on translated questions can enhance the cross-lingual transferability of reasoning tasks.
+There are also some works focusing on the cross-lingual transfer in the pre-LLM era.
+Devlin (2018) introduced mBERT, advancing cross-lingual transfer by capturing shared linguistic patterns in a unified embedding space, enabling zero-shot transfer without parallel corpora.
+Similarly, Conneau (2019) showed XLM’s effectiveness in optimizing multilingual embeddings, improving performance in translation and classification tasks.
+Ansell et al.
+(2021) proposed composable sparse fine-tuning, selectively fine-tuning sparse parameters across languages to reduce interference and boost performance, especially in low-resource settings, outperforming adapter-based methods in tasks like NER and NLI.
 Analysis of multilingual internal representation
@@
 The use of LLaMA-2 models, as well as several other large language models, complies with their respective licenses.
-Appendix A Detailed of Dataset
+Acknowledgements
+We would like to thank the anonymous reviewers for their insightful comments.
+Shujian Huang is the corresponding author.
+This work is supported by National Science Foundation of China (No.
+62376116, 62176120), research project of Nanjing University-China Mobile Joint Institute, the Fundamental Research Funds for the Central Universities (No.
+2024300507).
+Appendix A Details of Dataset
 A.1 Detailed description of Knowledge-Free Reasoning Dataset
@@
 We provide the results of the QASC dataset in Figure A3 and A4. The results show that the more knowledge provided leads to better cross-lingual transferability, which aligns with our conclusion.
-Appendix F Language Distribution of Model Training Corpora
-In this section, we present the language distribution of the pre-training corpora, referencing Table A6 from the LLaMA2 paper (Touvron et al., 2023) and Table A7 from the BLOOM paper (Workshop et al., 2022).
+Appendix F Evaluation on MMLU
+In order to address the concerns of overfitting on the fine-tuned datasets and forgetting the world knowledge, we also evaluate the LLaMA-2-7B-Chat model on MMLU before and after fine-tuning.
+The results are listed in Table A6.
+We find that the model’s MMLU performance shows both increases and decreases after finetuning, with relatively small changes in magnitude.
+This indicates that we do not cause the model to overfit these datasets and maintain its general capabilities.
+Finetuned Dataset MMLU(5-shot)
+None 40.88
+KFRD Arithmetic 45.53
+KFRD Symbolic 43.87
+KFRD Logical 45.14
+StrategyQA WF-all 46.85
+ASDiv 38.31
+Coin Flip 43.58
+ProofWriter 37.72
+Table A6: The accuracy of LLaMA-2-7B-Chat model on MMLU both before and after finetuning
+Appendix G Language Distribution of Model Training Corpora
+In this section, we discuss the language distribution of the pre-training corpora, from the LLaMA2 paper (Touvron et al., 2023) and the BLOOM paper (Workshop et al., 2022).
 Unfortunately, we were unable to locate the corresponding distribution data for Mistral and Qwen.
-For LLaMA2, languages such as Arabic and Hebrew were not included in the table, indicating that their proportions are lower than 0.005%, categorizing them as extremely low-resource languages.
+For LLaMA2, languages such as Arabic and Hebrew were not included in the table provided in Touvron et al.
+(2023) (Table 10), indicating that their proportions are lower than 0.005%, categorizing them as extremely low-resource languages.
 The other eight languages discussed in the paper are represented.
 Notably, German and Chinese rank as high-resource languages, accounting for 0.17% and 0.13% of the corpus, respectively, holding the second and fifth highest positions.
-For BLOOM, only English, French, Chinese, and Arabic are explicitly listed, while other languages are not reported in the table.
-Language Percent Language Percent
-en 89.70% uk 0.07%
-unknown 8.38% ko 0.06%
-de 0.17% ca 0.04%
-fr 0.16% sr 0.04%
-sv 0.15% id 0.03%
-zh 0.13% cs 0.03%
-es 0.13% fi 0.03%
-ru 0.13% hu 0.03%
-nl 0.12% no 0.03%
-it 0.11% ro 0.03%
-ja 0.10% bg 0.02%
-pl 0.09% da 0.02%
-pt 0.09% sl 0.01%
-vi 0.08% hr 0.01%
-Table A6: Language distribution in pretraining data with percentage >= 0.005%.
-Most data is in English, meaning that Llama 2 will perform best for English-language use cases.
-The large unknown category is partially made up of programming code data.
-Language Size in Bytes Language Size in Bytes
-Akan 70,1554 Arabic 74,854,900,600
-Assamese 291,522,098 Bambara 391,747
-Basque 2,360,470,848 Bengali 18,606,823,104
-Catalan 17,792,493,289 Chichewa 1,187,405
-chiShona 6,638,639 Chitumbuka 170,360
-English 484,953,009,124 Fon 2,478,546
-French 208,242,620,434 Gujarati 1,199,986,460
-Hindi 24,622,119,985 Igbo 14,078,521
-Indonesian 19,972,325,222 isiXhosa 14,304,074
-isiZulu 8,511,561 Kannada 2,098,453,560
-Kikuyu 359,615 Kinyarwanda 40,428,299
-Kirundi 3,272,550 Lingala 1,650,804
-Luganda 4,568,367 Malayalam 3,662,571,498
-Marathi 1,775,483,122 Nepali 2,551,307,393
-Northern Sotho 1,764,506 Odia 1,157,100,133
-Portuguese 79,277,543,375 Punjabi 1,572,109,752
-Sesotho 751,034 Setswana 1,502,200
-Simplified Chinese 261,019,433,892 Spanish 175,098,365,045
-Swahili 236,482,543 Tamil 7,989,206,220
-Telugu 29,934,071,59 Traditional Chinese 762,489,150
-Twi 1,265,041 Urdu 2,781,329,959
-Vietnamese 43,709,279,959 Wolof 3,606,973
-Xitsonga 707,634 Yoruba 89,695,835
-Programming Languages 174,700,245,772
-Table A7: Linguistic makeup of the ROOTS corpus. The ROOTS corpus is a multilingual dataset used for pretraining the BLOOM model.
-Appendix G Reasons for Creating a New Dataset
+For BLOOM, only English, French, Chinese, and Arabic are explicitly listed in the paper (Workshop et al., 2022, Table 1), while other languages are not reported in the paper, indicating they are low-resource languages.
+Appendix H Reasons for Creating a New Dataset
 The primary reason for creating a new dataset is that most existing datasets involve knowledge retrieval, which does not align with our focus on knowledge-free reasoning.
@@
 Table A7: Examples of adapted datasets used in this paper
-Appendix H Additional Works on Cross-lingual Transfer
-Devlin (2018) introduced mBERT, advancing cross-lingual transfer by capturing shared linguistic patterns in a unified embedding space, enabling zero-shot transfer without parallel corpora.
-Similarly, Conneau (2019) showed XLM’s effectiveness in optimizing multilingual embeddings, improving performance in translation and classification tasks.
-Ansell et al.
-(2021) proposed composable sparse fine-tuning, selectively fine-tuning sparse parameters across languages to reduce interference and boost performance, especially in low-resource settings, outperforming adapter-based methods in tasks like NER and NLI.
-Additionally, Ye et al.
-(2023) assessed the imbalance of knowledge across different languages in LLMs, observing weak cross-lingual transferability of knowledge.
-Zhu et al.
-(2024) discovered that training on translated questions can enhance the cross-lingual transferability of reasoning tasks.